### v1

# A Catalog of $\sim 146,000$ Radio Galaxies up to $z\sim 5$

Abstract
We present the results of optical identifications for Very Large Array Sky Survey (VLASS) radio sources employing the Ultraviolet Near Infrared Optical Northern Survey (UNIONS). A cross-match between UNIONS and VLASS Epoch 2 catalogs yields $146,212$ radio galaxies down to $r=24.5$ mag over a wide area of $\sim 4,200\mathrm{\ deg^{2}}$ in the northern hemisphere. We perform $g-$dropout selections and find $\gtrsim 200$ sources at $z\sim 4$ optimistically. Of 63,019 sources with valid photometric redshifts, 8,692 are at ${z}_{\mathrm{photo}}\geq 1$ and 1,171 are at ${z}_{\mathrm{photo}}\geq 2$. Based on spectral luminosities at 1.4 and 3 GHz using the valid UNIONS photo-z, we identify $\sim 49,000$ radio-loud AGNs (RLAGNs) with ${L}_{\mathrm{1.4GHz}}>10^{24}\mathrm{\ W\ Hz^{-1}}$, and all radio galaxies at $z\geq 1$ are RL. Adopting radio loudness instead, 138,266 out of 146,212 UNIONS-VLASS radio galaxies are RL. By virtue of the wide area and medium depth provided by UNIONS, our catalog greatly increases the number counts of RLAGNs at $z>1$. We further cross-match the UNIONS-VLASS catalog with LOw-Frequency ARray Two-metre Sky Survey (LoTSS) at 144 MHz and Faint Images of the Radio Sky at Twenty-cm (FIRST) at 1.4 GHz, yielding 101,671 UNIONS-VLASS-LoTSS, 79,638 -FIRST, and 64,672 -LoTSS-FIRST sources, respectively. This multifrequency radio dataset reveals sources of various spectral shapes, including the steep spectrum of aged populations, the peaked spectrum of young populations, and the upturned spectrum that might be associated with transient sources. The UNIONS-VLASS radio galaxy candidates will be covered by the Euclid wide survey, bringing about legacy values to benefit multi-faceted studies related to radio galaxies and beyond.
catalogs— galaxies: active — radio continuum: galaxies — surveys
facilities: VLA, CFHT, Pan-STARRS, Subaru HSC, and LOFARsoftware: astropy (Astropy Collaboration et al., 2013, 2018); astroquery (Ginsburg et al., 2019)
{CJK*}
UTF8gkai
show]yuxing.zhong.astro@gmail.com
show]ichikawa.waseda@gmail.com
1 Introduction
Supermassive black holes (SMBHs) are believed to ubiquitously reside in the center of most galaxies (Kormendy & Ho, 2013; McConnell & Ma, 2013), and a tight correlation between the mass of a SMBH and the properties of its host galaxy has been observed since the 1990s (Magorrian et al., 1998; Ferrarese & Merritt, 2000; Gebhardt et al., 2000; Tremaine et al., 2002). The tight correlation implies that SMBHs grow in tandem with the parameters of their host galaxies, so-called “co-evolution”. How SMBHs co-evolve with their host galaxies and how they impact their host galaxy evolution are long-standing questions in modern astronomy.
One key phenomenon to shape such co-evolution is called AGN feedback. Such a feedback can lead to the outflowing and/or heating of the cold gas, suppressing the host galaxy star formation and self-regulating the growth of the SMBH (Fabian, 2012; Alexander & Hickox, 2012; Bischetti et al., 2022; Harrison & Ramos Almeida, 2024; Talbot et al., 2024; Huško et al., 2025), although a compression of cold gas that enhances the star formation has also been observed (Shin et al., 2019; Duncan et al., 2023). One mechanism of AGN feedback is through the coupling between the host galaxy ISM and the relativistic plasma, which is known as radio jets, launched from the central SMBH. This radio/jet-mode feedback is considered the primary mode of AGN feedback in radio-loud active galactic nuclei (RLAGNs).
RLAGNs are often selected via the observed radio loudness $\mathcal{R}_{\mathrm{obs}}(={F}_{\mathrm{radio}}/{F}_{\mathrm{optical}}>10)$ defined as the ratio of radio-to-optical flux densities (Kellermann et al., 1989; Balokovic et al., 2012), or via the 1.4 GHz spectral luminosity ${L}_{\mathrm{1.4GHz}}\gtrsim 10^{24}\mathrm{\ W\ Hz^{-1}}$; Peacock et al. 1986; Condon 1989; Tadhunter 2016; Padovani et al. 2017). In the local Universe, RLAGNs are found to preferentially reside in early-type galaxies correlated with high stellar masses, low star formation rates, and radiatively inefficient AGNs (Best et al. 2005; Heckman & Best 2014; Gürkan et al. 2015; Pierce et al. 2022; however, see also Ramos Almeida et al. 2012; Pierce et al. 2023). At higher redshifts ($z>1$), there is an increasing number of RLAGNs found to be radiatively efficient (Eddington ratio ${\lambda}_{\mathrm{Edd}}\gtrsim 3\times 10^{-2}$; Merloni & Heinz, 2008; Ichikawa et al., 2021) and their host galaxies can be star-forming or even starburst galaxies (Drouart et al., 2014; Zhu et al., 2023; Zhong et al., 2024). These high-redshift radio galaxies (HzRGs) at $z>1$ are found tightly correlated with galaxy mergers (Chiaberge et al., 2015; Zhong et al., 2023; Heckman et al., 2024), suggestive of merger-triggered star formation and launching of radio jets. Therefore, RLAGNs across cosmic time, especially HzRGs, are valuable laboratories to study the interplay between the host galaxy evolution, SMBH growth, and radio jets, as well as the jet-accretion disc connections.
The rareness of RLAGNs ($10-20\%$ among all AGN populations; Kellermann et al. 1989; Urry & Padovani 1995; Ivezić et al. 2002; Chiaberge & Marconi 2011) requires large surveys to increase the sample size to statistically understand RLAGNs, their host galaxies, and the connections in between. Over the past decades, following the groundbreaking Cambridge Catalogue of Radio Sources (e.g., Kellermann et al., 1969), many radio sky surveys have been conducted centered around $\sim 1$ GHz, including the National Radio Astronomy Observatory (NRAO) Very Large Array (VLA) Sky Survey (NVSS; 1.4 GHz; Condon et al., 1998) and VLA Faint Images of the Radio Sky at Twenty-cm (FIRST; 1.4 GHz; Becker et al., 1995), and Australian Square Kilometre Array Pathfinder (ASKAP) Evolutionary Map of the Universe (EMU; 944 MHz; Norris et al., 2011) and Rapid ASKAP Continuum Survey (RACS; $\sim 900-1600$ MHz; McConnell et al., 2020). The vast sky coverages ($\mathrm{>10,000\ deg^{2}}$) of these surveys offer great opportunities to build up an unbiased sample to elevate the statistical importance of studies related to RLAGNs.
Wide-area optical surveys are necessary to identify the optical counterparts of radio sources and determine their redshifts, either photometrically or spectroscopically. Using the Sloan Digital Sky Survey (SDSS; York et al. 2000), Ivezić et al. (2002) and Helfand et al. (2015) identified the optical counterparts for the FIRST radio core sources at 1.4 GHz, adopting a positional cross-match between the two catalogs. Due to the shallow optical sensitivity of SDSS $r$-band imaging ($r<22.2$ mag), the fraction of matches (${f}_{\mathrm{matches}}$) is only $\sim 30\%$, and a significant fraction of optically faint host galaxies may have been missed. Using the Subaru Hyper Suprime-Cam (HSC) Subaru Strategic Program (SSP; Aihara et al., 2018), Yamashita et al. (2018) reaches ${f}_{\mathrm{matches}}=50.6\%$ optimistically for FIRST sources in the HSC field thanks to the depth of HSC reaching a limiting magnitude of 25.9 in i-band. Despite an improvement in associating radio sources with faint objects, elevating the fairness of the radio galaxy catalog, the survey area is $\mathrm{1,400\ deg^{2}}$ such that the catalog size is limited.
Other than the 1.4 GHz continuum, radio galaxy catalogs have been constructed centering at 3 GHz and 144 MHz. Based on the largest sample of uniformly selected and spectrally confirmed quasars provided by SDSS, Arnaudova et al. (2024) identified 3,697 RL and 111,132 RQ sources in LOw-Frequency ARray (LOFAR) observations at 144 MHz via the radio loudness dichotomy. With 384 h observations of the upgraded Karl G. Jansky VLA, the VLA-COSMOS 3 GHz Large Project (Smolčić et al., 2017) identifies 10,830 radio sources in the COSMOS field. This project is equipped with a sensitivity better than that of FIRST by almost two orders of magnitude, greatly exploring the RQ populations. These catalogs are advantageous explorations in the faint end of radio galaxies; however, they do not significantly expand the number counts of RLAGNs.
To improve both the sample size of RLAGNs and the unbiasedness of host galaxy populations, we search for UNIONS (Gwyn et al., 2025) optical counterparts for VLASS (Lacy et al., 2020) radio sources at 3 GHz. UNIONS reaches a depth of 24.7 mag in r-band and covers an extremely wide area of $\sim 5,000\mathrm{\ deg^{2}}$ that fully overlaps with VLASS, bridging the gap between the shallow imaging of SDSS and the deep but slightly limited sky coverage of HSC-SSP. VLASS is the state-of-the-art radio survey operated at 3 GHz in terms of spatial resolution, sensitivity, and total survey area. The cross-match between these two catalogs will produce one of the largest RLAGN catalogs covering a wide redshift range. Furthermore, both catalogs are ongoing projects, ensuring a great scalability of the UNIONS-VLASS catalog.
In this paper, we use radio galaxies for sources cross-matched between the UNIONS optical and VLASS radio catalogs, and define HzRGs as radio galaxies at $z>1$. This paper is organized as follows. In Section 2, we introduce the optical and radio source catalogs used to build the radio galaxy catalog. The cross-matches between the catalogs and the statistical results of our radio galaxy catalog are presented in Section 3. We examined the optical and radio number counts of UNIONS-VLASS radio galaxies in Section 4, and we further presented the selection of RLAGNs based on the observed radio loudness and spectral luminosity. We show radio galaxy classifications using multifrequency radio datasets and discuss the high-z sources based on the dropout technique in Section 5. A summary is presented in Section 6. Throughout this paper, we assume a $\mathrm{\Lambda CDM}$ cosmology with $\Omega_{m}=0.309$, $\Omega_{\Lambda}=0.691$, and $H_{0}=\mathrm{67.7\ km\ s^{-1}\ Mpc^{-1}}$ (Planck Collaboration et al., 2016).
2 The Data
This exploration for radio galaxies is established on the basis of angular separation-based cross-match between the VLASS for radio sources at 3 GHz and UNIONS for optical counterparts. The cross-matched UNIONS-VLASS catalog is set as our reference and further cross-matched with the FIRST 1.4 GHz and LoTSS 144 MHz radio source catalogs. In Figure 1, we summarize the flowchart that illustrates the criteria applied to each survey to filter out valid sources for the cross-match and to exclude invalid sources in the resultant cross-matches.
2.1 UNIONS
An elevation of the statistical importance of our radio galaxies catalog requires a wide ($>1000$ deg2) and deep ($r>22$ mag) optical survey. The Ultraviolet Near-Infrared Optical Northern Survey (UNIONS; Gwyn et al. 2025), as a science collaboration between wide imaging surveys that cover the northern hemisphere, provides an excellent opportunity to abundantly discover the optical counterparts of radio sources. The UNIONS integrates multiband (five broadbands, $u$, $g$, $r$, $i$, $z$) observations from Canada-France-Hawaii Telescope/Canada-France Imaging Survey (CFHT/CFIS, $u+r$), Subaru/Waterloo-Hawaii-IfA G-band Survey (WHIGS, g), The Panoramic Survey Telescope & Rapid Response System (Pan-STARRS, i), and Subaru/Wide Imaging with Subaru HSC of the Euclid Sky (WISHES, z). The main UNIONS footprint is located at a declination (DEC.) $>30^{\circ}$ and a galactic latitude of $|b|>25^{\circ}$, enclosing a total area of $4,861\mathrm{\ deg^{2}}$. Extended $u-$ and $z-$ (WISHES+) observations are being conducted to cover the sky area at $\mathrm{DEC.}<30^{\circ}$, elevating the sky coverage of UNIONS. Albeit UNIONS is motivated by the Euclid wide survey (Euclid Collaboration et al., 2022), efforts have been devoted to maintaining UNIONS as an independent survey with scientific values beyond what will be brought about by the Euclid.
The $r-$band observation of CFHT/CFIS has an excellent median seeing of $0.69\arcsec$ and reaches a surface brightness limit of 28.4 $\mathrm{mag/arcsec^{2}}$ for extended sources. It is set as the reference catalog by which the elliptical aperture is defined for the forced photometry applied to all five UNIONS bands. Since the broadband observations of UNIONS are conducted under multiple telescopes, the Gaussian Aperture and Photometry ($\mathtt{GAaP}$; Kuijken et al. 2015) is adopted to produce Gaussianized images for raw imaging data collected from observations at different instruments. The forced photometry with Gaussian-weighted, elliptical apertures can then be properly used to perform homogeneous multiband photometric measurements simultaneously. The depth of UNIONS is characterized by the $10\sigma$ point source depth measured through a $2.0\arcsec$ aperture. The current median depth of UNIONS is $[u,g,r,i,{z}_{\mathrm{PS}},{z}_{\mathrm{HSC}}]=[23.7,24.5,24.2,23.8,23.3,23.3]$, to be compared with the original goal of $[u,g,r,i,z]=[23.6,24.5,23.9,23.6,23.4]$ (Gwyn et al., 2025).
[…]
Figure 1: Flowchart of the cross-match processes and sample selections (see § 3 for details). In cyan boxes that describe the sample sizes, we supplement the sections where the corresponding sample is described and discussed.
Figure 2: The angular separation-$\mathtt{NN\_dist}$ plane for UNIONS-VLASS sources cross-matched with a search radius of $6^{\prime\prime}$. The vertical lines indicate the $\mathtt{NN\_dist}$ cut of $30^{\prime\prime}$ and the horizontal line represents search radius to estimate completeness and contamination.
2.2 VLASS
The VLASS (Very Large Array Sky Survey; Lacy et al. 2020) observes the entire sky at a declination $>-40$ degrees, covering a total area of 33,885 $\mathrm{deg^{2}}$. The observation frequency ranges $2-4$ GHz, with a central (reference) frequency of 3 GHz. In baseline configurations in B and BnA to compensate for the projection effect, an angular resolution of $2.5\arcsec$ has been reached. The total survey is divided into three epochs, which have been completed in October 2024, plus a further half-epoch observation. [^1] https://science.nrao.edu/vlass/library/vlass-epoch-4-science-case Plus the single-epoch rms sensitivity (root mean square error $1\sigma$) that reaches 120 $\mu$Jy, these features make VLASS the widest and most sensitive radio sky survey with an excellent resolution at 3 GHz (de Gasperin et al., 2021). Therefore, VLASS is ideal for serving as our reference radio source catalog.
VLASS data release delivers quick look (QL) and single epoch (SE) images. QL images are products of the rapid-CLEAN process with a cell size of $1\arcsec$. SE images undergo deeper cleaning and self-calibration and reach a better cell size of $0.6\arcsec$. However, only an area of 1,000 $\mathrm{deg^{2}}$ is available in SE, we therefore only use QL products for the cross-match. We use the epoch 2 catalog (VLASS2 hereafter) since astrometric errors and underestimates of flux densities were previously present in epoch 1, and the final release of epoch 3 is unsettled.
There are $\sim 3$ M radio sources detected in VLASS2. Prior to cross-matching, we adopted the following criteria to filter the VLASS2 catalog following the VLASS catalog user guide (see also Gordon et al. 2021)
  • $\mathtt{E\_RA\&E\_DEC<2.5^{\prime\prime}}$: astrometric uncertainties in both RA and DEC should be smaller than the angular resolution of $2.5\arcsec$;
  • $\mathtt{Duplicate\_flag<2}$: the source has no duplicated or has a preferred version of the duplicated counterpart within $2\arcsec$ search radius;
  • $\mathtt{Quality\_flag==(0|4)}$: the source is unflagged or its peak flux density is higher than the integrated one ($\mathtt{Peak\_flux>Total\_flux}$);
  • $\mathtt{S\_code!=E}$: the source can be fitted with a single-Gaussian or multi-Gaussian components;
plus $\mathtt{NN\_dist}$, which the angular separation between the target source and its nearest neighbor in VLASS2, is larger than $30^{\prime\prime}$ (see § 3.1 for the discussion). These criteria reduce the number of sources in “clean” VLASS2 to 1,380,707. The VLASS2 catalog also includes the column $\mathtt{P\_sidelobe}$, which means the probability of a source being a sidelobe but misclassified as a real source. As recommended in the VLASS Catalog User Guide, $\mathtt{Quality\_flag==0}$ that indicates $\mathtt{P\_sidelobe<0.1}$ is preferred, and the estimated contamination rate is 0.39%. In the cleanest criterion, $\mathtt{P\_sidelobe<0.05}$ is applied, and the estimated contamination rate is 0.12%. For the final “clean” sample of 146,212 UNIONS-VLASS sources cross-matched (see § 3.3 for details), adopting these two criteria results in a difference of 1,766 ($\sim 1.2\%$) sources, which is statistically insignificant. In the following of this work, without the loss of totality of our catalog, we present results adopting $\mathtt{P\_sidelobe<0.1}$.
2.3 FIRST
FIRST (Becker et al., 1995) employs VLA to observe 8,444 $\mathrm{deg^{2}}$ square degrees of the NGC at 1.4 GHz and the survey area is limited to a declination of $<65$ degrees. In the SGC, an area of $644\mathrm{\ deg^{2}}$ has been observed, whereas the survey area does not overlap with the current internal data release of UNIONS5000 (see Figure 5). In this work, we use the FIRST final data release catalog (Helfand et al., 2015). In B-configuration, the survey reaches a spatial resolution represented by a circular Gaussian with full width at half maximum (FWHM) $5.4\arcsec$ with a pixel size of $1.8\arcsec$ over the northern sky. For bright, pointlike sources with measured extents of $\leq 2.0\arcsec$, the statistical astrometry is $0.22\arcsec$. In a typical situation, the flux density detection limit is 1 mJy over the full footprint.
2.4 LoTSS
To understand the evolution of radio galaxies, the spectral shape plays a crucial role, which requires multifrequency observations to construct the synchrotron spectral energy distribution (SED). MHz-observations carried out by LOw-Frequency ARray are ideal supplements to our GHz-VLASS and FIRST data. The LoTSS (LOFAR Two-metre Sky Survey; Shimwell et al. 2017, 2022) employs the LOFAR to observe 27% of the northern sky at $120-168$ MHz, with a central (reference) frequency of 144 MHz. The sky coverage has two separate regions that span 4,178 and 1,457 $\mathrm{deg^{2}}$ respectively. The full bandwidth Stokes I continuum imaging of LoTSS yields images at a spatial resolution of $6.0\arcsec$, centered at a frequency of 144 MHz. The median detection sensitivity of the survey is 89 $\mathrm{\mu Jy\ beam^{-1}}$ with a typical positional uncertainty smaller than $0.2\arcsec$. In this work, we use the latest and widest coverage catalog, LoTSS DR2 (Shimwell et al., 2022), which contains 4,396,228 radio sources derived from integrated intensity maps in Stokes I polarization.
3 UNIONS-VLASS Radio Galaxy Catalog
In this section, we describe how we construct the UNIONS-VLASS radio galaxy catalog and how we supplement this catalog with multifrequency radio surveys, including FIRST and LoTSS. We also describe the key information to be presented in our catalog, including redshifts, luminosities, spectral indices, and how a radio source is classified as an extended source.
Figure 3: Leftmost: Tile-wise number densities ${n}_{\mathrm{den}}$ of galaxies per $\mathrm{arcmin^{2}}$ for UNIONS tiles in the North Galactic Cap. The violet, blue, and orange lines indicate average, median, and maximal ${n}_{\mathrm{den}}$, respectively. Middle: completeness and cleanness as a function of the search radius ${r}_{\mathrm{s}}$. The vertical dashed lines mark where the cleanness approximately intersects the completeness, indicating a trade-off between the totality of the catalog and the exclusion of the contamination. Rightmost: differential distribution of separation between UNIONS optical and VLASS radio sources including radio sources that are associated with multiple optical counterparts.
3.1 $\mathtt{NN\_dist}$ Cut
Radio galaxies with low radio luminosities often have their jets confined by the ISM and rapidly disrupted (Fanaroff & Riley, 1974; Turner & Shabala, 2015). As a result, the peak surface brightness can be close to the host galaxy. The nearest neighbour matching is thus a straightforward and efficient approach for identifying the optical counterpart of a radio source (Williams et al., 2019). On the other hand, powerful radio jets can rapidly overcome the confinement of the ISM of their host galaxies without suffering significant disruption (Mukherjee et al., 2016). Thereby, they often propagate far from their host galaxies and end in the terminal shocks known as radio hotspots, where the observed radio surface brightness peaks (Turner & Shabala, 2023). Identifying the optical counterparts of such radio sources requires a larger search radius, which, in turn, inevitably increases the likelihood of contamination by unrelated sources (Banfield et al., 2015; Gordon et al., 2023).
To statistically avoid erroneous associations attributed to large-scale, complex radio structures, we perform a cross-match between UNIONS and the VLASS2 catalog – adopting the criteria present in Gordon et al. (2021) – choosing a search radius of $6^{\prime\prime}$, which is $\sim 550,000$ sources. The angular separation between UNIONS-VLASS versus $\mathtt{NN\_dist}$ is shown in Figure 2. It is clear that the number of cross-matched sources shows a bimodal distribution along $\mathtt{NN\_dist}$. Sources cross-matched with $\mathtt{NN\_dist<30^{\prime\prime}}$ tend to have large angular separations between UNIONS and VLASS centroids. We randomly select 36 VLASS sources with $\mathtt{NN\_dist<10^{\prime\prime}}$ and show their QL images, overlaid with VLASS and UNIONS source centroids, in Figure 18. For these radio sources, they often show complex and extended radio morphology, including core-jet and double-jet structures. VLASS could capture the centroid of the lobe or radio hotspot, naturally begetting the large offsets between UNIONS and VLASS. We, therefore, choose a $\mathtt{NN\_dist}$ cut of $30^{\prime\prime}$ to avoid false positive associations between UNIONS-VLASS arising from evolved radio structures that can extend well beyond host galaxy scale and overlap with other galaxies. This cut can remove a large fraction of bright, evolved radio galaxies at low-$z$. Consequently, we see that UNIONS-VLASS RGs are absent from the bright end (${F}_{\mathrm{3GHz}}>3000$ mJy; Figure 9) and could be biased toward compact, young radio AGN populations (Figures 1516; see § 5.3 for discussion).
3.2 Contamination and Completeness
Since the galaxies are randomly distributed throughout the whole sky, the one-to-one correspondence between the radio and optical sources could be a false positive. To statistically circumvent such contamination, we estimate the cleanness and completeness of the catalog in the following way.
We first estimate the number density of galaxies in each UNIONS tile lying within the VLASS field. This gives a maximum number density of ${n}_{\mathrm{den,max}}=81.1$ per $\mathrm{arcmin^{2}}$ and a median value of ${n}_{\mathrm{den,med}}=36.1$ per $\mathrm{arcmin^{2}}$ (leftmost panel of Figure 3). We perform an angular separation-based cross-match, choosing a search radius of $2.5\arcsec$. We remove optical counterparts without valid photometry in any of ugriz and multiple optical counterparts corresponding to a single radio source. The total number of clean UNIONS-VLASS radio-optical pairs is then ${N}_{\mathrm{tot}}$, and the separation between each pair is defined as ${r}_{\mathrm{sep}}$. Within the search radius of $r_{s}$, the number of sources with ${r}_{\mathrm{sep}}\leq{r}_{\mathrm{s}}$ is ${N}_{\mathrm{s}}$, and these sources occupy a total area of ${A}_{\mathrm{s}}=\pi{r}_{\mathrm{s}}^{2}{N}_{\mathrm{s}}$ $\mathrm{arcsec^{2}}$. The completeness of UNIONS-VLASS is then represented by ${N}_{\mathrm{s}}/{N}_{\mathrm{tot}}$ and the maximum is 100%. The number of optical sources adjacent to radio sources by coincidence is ${N}_{\mathrm{rand}}={A}_{\mathrm{s}}{n}_{\mathrm{den}}$. The cleanness is then calculated by $({N}_{\mathrm{tot}}-{N}_{\mathrm{rand}})/{N}_{\mathrm{tot}}$. The ${r}_{\mathrm{s}}$ at which the cleanness intersects the completeness represents a compromise between the totality of the catalog and the contamination control (middle panel of Figure 3). With ${n}_{\mathrm{den,med}}=36.1$ per $\mathrm{arcmin^{2}}$, the contaminated cross-matches arising from chance coincidence is 7.5%, approaching an estimated completeness of 92.7% at ${r}_{\mathrm{s}}\sim 1.6\arcsec$. When ${n}_{\mathrm{den,max}}=81.1$ per $\mathrm{arcmin^{2}}$ is adopted, ${r}_{\mathrm{s}}\sim 1.2\arcsec$ is preferred, and the contamination rate is 9.1% while the totality of the catalog is 89.3%.
Without excluding optical counterparts associated with a single radio source, we show the differential distribution of the separation between UNIONS and VLASS sources (rightmost panel of Figure 3). At ${r}_{\mathrm{s}}>1.5\arcsec$, the increase in the search radius is accompanied by an increase in the number of optical counterpart counts in the annulus, suggesting these optical counterparts are likely to be coincident associations rather than real matches. The threshold of ${r}_{\mathrm{s}}\sim 1.5\arcsec$ is consistent with the completeness-cleanness trade-off of $1.6\arcsec$ adopting a median galaxy number density of $36.1$ per $\mathrm{arcmin^{2}}$. This value is also comparable to Gordon et al. (2023), who found that the probability that an AllWISE soruce is the true positive counterpart of an unresolved VLASS radio source reaches 80% at a positional uncertainty of $1.8\arcsec$.
Table 1: Summary of the cross-match results
${N}_{\mathrm{VLASS2}}$ ${N}_{\mathrm{UNIONS}}$ ${N}_{\mathrm{matches}}$ ${f}_{\mathrm{matches}}$ Area
[$\mathrm{deg^{2}}$]
246,325 670,038,362 146,212 0.594 $\sim 4,200$
From left to right, the column represents number of VLASS2 sources, number of UNIONS sources, number of matches, fraction of matches, and the total cross-matched area.
3.3 Clean UNIONS-VLASS Radio Galaxies
Choosing an angular separation of $\Delta\theta=\sqrt{(\Delta\alpha\cos\delta)^{2}+(\Delta\delta)^{2}}\leq 1.5\arcsec$ as the search radius, our cross-match between the UNIONS and VLASS Epoch 2 catalogs immediately yields 164,169 optical counterparts of radio sources. This search radius represents a trade-off between the completeness of the catalog and a statistical exclusion of optical sources coincidentally associated with radio sources, which is described in § 3.2.
To further make a clean radio galaxy catalog, we first eliminate optical sources without valid photometry in $r-$band since the UNIONS $\mathtt{GAaP}$ catalog is constructed upon the successful CFHT $r-$band detections. This reduces the number of optical sources to 157,346. Out of these sources, there are 135,671 UNIONS sources uniquely associated with VLASS sources, and they are flagged with $\mathtt{0}$ in the column $\mathtt{flag\_multi\_opt}$.
If the radio source has more than one optical counterpart, we always choose the nearest one as the real match. This yields 10,541 radio sources that have multiple optical sources within a search radius of $1.5\arcsec$, and they are flagged with $\mathtt{1}$ in the column $\mathtt{flag\_multi\_opt}$. Because of the general trend of being compact for our UNIONS-VLASS RGs (see § 4.1 and § 5.1), this nearest match method could be valid in most cases, as jets and jet-induced shocks have not traveled far away from the host galaxy. However, it should be noted that this does not fully rule out optical counterparts by chance overlaps. Especially in cases where the RGs have developed jet-lobe structures that can result in large offsets between radio-emitting regions and host galaxies. Consequently, the centroid of the radio emission captured by VLASS may be the radio hotspot and appears to be closer to another galaxy (up to approximately $4\%$ based on $\sim 1,500$ Euclid-VLASS RGs; Zhong et al. in prep.). These multiple optical counterparts are also considered merging or interacting galaxy candidates, and they will be investigated in detail in future work.
The results of the cross-match are summarized in Table 1. Our angular separation-based cross-match generated 146,212 VLASS radio sources matched to UNIONS optical counterparts. Defining the fraction of matched VLASS sources in the area that overlaps with the UNIONS as the fraction of matches $({f}_{\mathrm{matches}})$, the corresponding value is 0.594, which is the maximum to be considered. By virtue of the higher sensitivity and wide sky coverage provided by the UNIONS, this rate is higher than those in the previous studies cross-matching FIRST with SDSS ($\sim 30\%$; Ivezić et al. 2002; Helfand et al. 2015). This fraction of matches also exceeds the cross-matching FIRST with Subaru HSC ($\sim 47\%$; Yamashita et al. 2018), even HSC optical number counts share a similar trend as the UNIONS (see 4.2 for further discussions). This could be explained by a combination of several factors, including a higher sensitivity of VLASS over FIRST and a larger search radius ($1.5\arcsec$) of UNIONS-VLASS than $1\arcsec$ of HSC-FIRST.
In summary, our cross-match between UNIONS and VLASS catalogs yields 146,212 radio galaxies. This “clean” UNIONS-VLASS radio galaxy catalog is the final sample to be presented in this work.
Figure 4: Astrometric offsets in R.A. ($\Delta\alpha$) and DEC. ($\Delta\delta$) for 146,212 “clean” UNIONS-VLASS sources (blue) associated via $\sqrt{\Delta\alpha^{2}+\Delta\delta^{2}}\leq 1.5\arcsec$ and subsample of 140,065 source adopting positional cross-match (orange; see texts for details). Mean values of the offsets are embedded in the bracket and $\sigma$ represents the standard deviation. The typical astrometric offset of VLASS Epoch 1 is $\sim 0.5\arcsec$ at $\mathrm{DEC.>-20\ deg^{2}}$.
3.4 UNIONS-VLASS Astrometry
This vast radio galaxy catalog offers an excellent opportunity to investigate the systematic astrometric offsets between UNIONS and VLASS catalogs. The offsets along R.A. and DEC. are residuals of a VLASS position minus a UNIONS position: $\Delta\alpha=\mathrm{\alpha_{VLASS}-\alpha_{UNIONS}}$ and $\Delta\delta=\mathrm{\delta_{VLASS}-\delta_{UNIONS}}$. However, our cross-match method selects associated pairs based on the angular separation ($\Delta\theta$) between radio and optical sources. Because the positional difference along the R.A. on the celestial sphere has a declination dependence, this calculation has to be corrected for the projection effect and is given by $\Delta\theta=\sqrt{({\Delta\alpha}_{\mathrm{corr}})^{2}+(\Delta\delta)^{2}}$, ${\Delta\alpha}_{\mathrm{corr}}=(\mathrm{\alpha_{VLASS}-\alpha_{UNIONS}})\times\cos({\delta}_{\mathrm{UNIONS}})$. Therefore, for $\Delta\alpha=\mathrm{\alpha_{VLASS}-\alpha_{UNIONS}}$, the value can exceed the search radius of $1.5\arcsec$, and such an excess becomes more severe for sources at high declinations, as shown in the left panel of Figure 19.
To make a fair comparison of the astrometric accuracy in R.A. with other catalogs adopting positional cross-match that simply calculates the residuals between coordinates, we then make a subsample of “clean” UNIONS-VLASS sources. This subsample contains 140,065 UNIONS-VLASS sources that have their optical counterparts lying within a circle of a projected radius $1.5\arcsec$ on the sky and centered on the radio source. This reduces the catalog size by $\sim 4.2\%$, which is statistically insignificant. The distributions of $\Delta\alpha$ and $\Delta\delta$ of these sources are shown by orange histograms in Figure 4.
The standard deviations of 140,065 UNIONS-VLASS sources whose $\sqrt{\Delta\alpha^{2}+\Delta\delta^{2}}\leq 1.5\arcsec$ are $0.44\arcsec$ in R.A. and $0.29\arcsec$ in DEC. On the other hand, for 146,212 “clean” UNIONS-VLASS sources, the offset along the R.A. retards to $0.61\arcsec$ while the change in the offset along DEC. is negligible. These offsets are much worse than the positional uncertainties (20 mas) of the CFHT $u-$ and $r-$band imaging surveys employed by UNIONS (Gwyn et al., 2025). For the VLASS catalog, by comparing the Epoch 1 catalog with Gaia DR2, the VLASS component positional accuracy is $\sim 0.5\arcsec$ at $\mathrm{DEC.>-20\ deg^{2}}$. Therefore, the current standard deviations of the positional errors for our UNIONS-VLASS sources are thus reasonable compared to the entire VLASS survey, suggesting that the astrometric offsets between optical and radio pairs are attributed to the intrinsic uncertainties of VLASS QL images.
In VLASS Epoch 1, the positional error peaks at $0.0\arcsec$ in R.A. and peaks at $-0.25\arcsec$ in DEC. Our UNIONS-VLASS sources based on Epoch 2 observations show a similar trend in R.A. with an improved astrometric accuracy in DEC. (${\Delta\delta}_{\mathrm{peak}}\sim 0.0\arcsec$) compared to VLASS Epoch 1. However, there is no significant improvement in the typical astrometry of $\sim 0.5\arcsec$, which is much worse than that of FIRST ($0.2\arcsec$). This is more of a problem of the rapid CLEAN procedure adopted by VLASS QL products. More accurate astrometry can be expected when Epoch 3 and epoch-concatenated data are delivered, as well as for the SE images.
3.5 Cross-matches between Surveys
Once the unique optical counterpart catalog of VLASS sources (146,212 “UNIONS-VLASS Radio Galaxies” in Figure 1) is constructed, we then perform additional cross-match with the VLA/FIRST 1.4 GHz survey and LOFAR/LoTSS 144 MHz survey. We search for FIRST 1.4 GHz radio counterparts around our UNIONS-VLASS radio galaxies by requiring the angular separation between a pair of VLASS and FIRST sources to be within a $2.5\arcsec$ radius. This results in a total of 85,060 matched sources. We then remove the sources containing possible sidelobes, using the sidelobe probability $\mathtt{P(S)}$ presented in the FIRST catalog. We choose $\mathtt{0.00<P(S)<0.05}$ as the threshold, by which 76.3% of sources are considered true positives, and this results in 79,638 radio sources identified in UNIONS, VLASS, and FIRST.
We search for LoTSS 144 MHz radio counterparts for the UNIONS-VLASS radio galaxies by applying a positional cross-match of $<2.5\arcsec$ radius, consistent with the VLASS beam size. This cross-match results in 101,671 radio galaxies identified in UNIONS, VLASS, and LoTSS. Finally, there are 64,672 UNIONS-VLASS-FIRST-LoTSS radio galaxies.
3.6 Catalog Information
Figure 5: The sky coverages of UNIONS (black), VLASS (cyan), FIRST (green), and LoTSS (orange).
In Figure 5, we show the sky coverages of UNIONS (black), VLASS (cyan), FIRST (green), and LoTSS (orange). The presented “clean” UNIONS-VLASS catalog contains 146,212 radio galaxies that are distributed amongst an area of $\sim 4,200\mathrm{\ deg^{2}}$, reaching a number density of $\sim 34\ \mathrm{deg^{-2}}$. In the following subsections, we introduce the key information – as well as how they are computed – to be included in the catalog.
3.6.1 Valid Photometric Redshifts of UNIONS
The photometric redshifts presented in UNIONS are estimated from Bayesian Photometric Redshift (BPZ; Benítez, 2000). The source with reliable photometric redshift should have valid photometric measurements in all five bands ($ugriz$): (1) the photometry is not flagged (column $\mathtt{FLAG\_GAaP\_{ugriz}==0}$) and (2) the AB magnitudes within the optimal Gaussianized aperture should be valid measurements (column $\mathtt{MAG\_GAaP\_{ugriz}!=(-99|99})$), where $\mathtt{99}$ indicates that the source is below the detection limit and $\mathtt{-99}$ stands for an absence of the observational data. Therefore, of 146,212 “clean” UNIONS-VLASS sources, 20,358 have photo-$z$ with good reliabilities. These sources are labeled with $\mathtt{0}$ in the column $\mathtt{flag\_z\_photo}$.
[…]
3.6.2 Photometric versus Spectroscopic Redshifts
Gwyn et al. (2025) compared ${z}_{\mathrm{photo}}$ estimates with the spectroscopic redshifts collected from publicly available datasets in the UNIONS footprint. The bias and scatter of photo-$z$ are represented by the median and normalized median absolute deviation (NMAD) of the quantity $\frac{\Delta z}{1+z}=\frac{{z}_{\mathrm{spec}}-{z}_{\mathrm{photo}}}{1+{z}_{\mathrm{spec}}}$, respectively. Correspondingly, the bias is $-0.008$ and the scatter is $0.061$ for UNIONS sources with ${z}_{\mathrm{photo}}\lesssim 1.6$. If selecting outliers by $\frac{|\Delta z|}{1+z}>0.15$, then $16.3\%$ of the sources are considered having no reliable photo-$z$ estimates.
Figure 6: Comparisons between UNIONS photometric and SDSS spectroscopic redshifts for ${\color[rgb]{0,0,0}2,175}$ UNIONS-VLASS sources. The solid line indicates ${z}_{\mathrm{photo}}={z}_{\mathrm{spec}}$ while the dashed lines indicate $({z}_{\mathrm{photo}}-{z}_{\mathrm{spec}})/(1+{z}_{\mathrm{spec}})=\pm 0.1$, which represents the photo-$z$ scatter for guidance.
[…]
Figure 7: The ratio of the total flux density to the peak flux density as a function of S/N defined as the ratio of the peak flux density to the survey detection limit for VLASS (blue crosses) and FIRST (orange pluses). The dashed and solid lines represent the lower and upper envelopes, respectively (see text for details). Those lying above the upper envelope are considered extended sources. The plus and crosses represent the examples of extended and compact sources, respectively, shown in Figure 8
Figure 8: We show examples of extended (upper) and compact sources (lower) selected from the criterion described in § 3.6.3 to demonstrate the validity. The corresponding positions of these two sources ${F}_{\mathrm{int}}/{F}_{\mathrm{peak}}$ and S/N are shown in Figure 7. The white ellipses in the bottom left corner of each panel indicate the beam size. For both rows, from left to right, the cutout is the product of VLASS Epoch 2 QL, Epoch 3 QL, and SE image, respectively.
3.6.3 Extended Radio Sources
The morphology of radio emission, such as its compactness, presumably originating from radio jets, provides key information about the evolution of RLAGNs. To distinguish between compact (unresolved) and extended (resolved) sources, we make use of the ratio of the total radio flux density (${F}_{\mathrm{int}}$) to the peak radio flux density (${F}_{\mathrm{peak}}$), a widely adopted method (e.g. Ivezić et al., 2002; Bondi et al., 2003; Schinnerer et al., 2007). This ratio directly measures the extension of a radio source, and for unresolved or marginally resolved sources, ${F}_{\mathrm{int}}={F}_{\mathrm{peak}}$ (e.g., Bondi et al., 2008). Extended and resolved sources have ${F}_{\mathrm{int}}>{F}_{\mathrm{peak}}$ whereas the background noise could lower the total flux density. Therefore, we plot ${F}_{\mathrm{int}}/{F}_{\mathrm{peak}}$ as a function of the signal-to-noise ratio $\mathrm{S/N}\ (={F}_{\mathrm{peak}}/\mathrm{rms})$ in Figure 7 for 146,212 “clean” UNIONS-VLASS sources (blue crosses) and those (79,638) with FIRST detections (orange dots). Values of ${F}_{\mathrm{int}}<{F}_{\mathrm{peak}}$ then a result of the measurement uncertainties probably arising from the background noises that could lower ${F}_{\mathrm{int}}$ (e.g., Bondi et al., 2003).
To select extended sources, we first determined a lower envelope that encloses 99% of the UNIONS-VLASS sources with ${F}_{\mathrm{int}}<{F}_{\mathrm{peak}}$, as indicated by the dashed cyan curve. We then mirrored it above the horizontal line defined by ${F}_{\mathrm{int}}/{F}_{\mathrm{peak}}=1$, and this upper envelope (solid cyan curve) is represented as:
${F}_{\mathrm{int}}/{F}_{\mathrm{peak}}=1+8.6\times\left({{F}_{\mathrm{peak}}/\mathrm{rms}}\right)^{-1}.$ (1)
Adopting these criteria, 29,862 ($\sim 20\%$) of our UNIONS-VLASS sources are extended. We performed a similar evaluation for sources detected in FIRST where the FIRST S/N is defined as $({F}_{\mathrm{peak}}-0.25)/\rm{rms}$ to correct for the CLEAN bias (Helfand et al., 2015), and the corresponding upper envelope is defined by ${F}_{\mathrm{int}}/{F}_{\mathrm{peak}}=1+7.9\times\left[\left({{F}_{\mathrm{peak}}-0.25}\right)/\mathrm{rms}\right]^{-1}$, which is in general agreement with ${F}_{\mathrm{int}}/{F}_{\mathrm{peak}}=1+6.5\times\left({{F}_{\mathrm{peak}}/\mathrm{rms}}\right)^{-1}$ calculated for the HSC-FIRST catalog (Yamashita et al., 2018).
Alternatively, thanks to the higher sensitivity and better resolution, the VLA-COSMOS 3 GHz survey (Smolčić et al., 2017) determined the envelope by a 95% enclosure of the sources below the ${F}_{\mathrm{int}}={F}_{\mathrm{peak}}$ line. Adopting this criterion, the upper envelope is
${F}_{\mathrm{int}}/{F}_{\mathrm{peak}}=1+3.7\times\left({{F}_{\mathrm{peak}}/\mathrm{rms}}\right)^{-1},$ (2)
and 63,203 ($\sim 43\%$) of our UNIONS-VLASS sources are extended. In this work, we adopt Eq. (1) as the threshold considering the similar beam sizes and sensitivities between VLASS and FIRST, and extended radio sources are labeled with $\mathtt{1}$ in column $\mathtt{extended}$. Examples of extended and compact radio sources are shown in Figure 8 and they are marked on the ${F}_{\mathrm{int}}/{F}_{\mathrm{peak}}-\mathrm{S/N}$ plane (Figure 7).
3.6.4 Additional Columns
[…]
We calculate the angular separations between each pair of UNIONS, VLASS, FIRST, and LoTSS sources, and they are recorded in the corresponding columns starting with $\mathtt{sep\_\{survey1\}\_\{survey2\}}$. We calculate the spectral indices between each pair of radio surveys if there are detections, and the corresponding columns start with $\mathtt{alpha\_\{frequency1\}\_\{frequency2\}}$. Since the spectral index is sensitive to uncertainties in the flux densities, we provide the corresponding uncertainties in the spectral index in the columns $\mathtt{alpha\_\{frequency1\}\_\{frequency2\}\_err}$.
We calculate the spectral luminosities [$\mathrm{\ W\ Hz^{-1}}$] at 1.4 and 3 GHz using the following equation:
$L_{\nu}=\frac{S_{\mathrm{obs}}4\pi D^{2}_{L}}{(1+z)^{1+\alpha}},$ (3)
[…]
  • detected in FIRST but not in LoTSS: $S_{\mathrm{obs}}=S_{\mathrm{FIRST}}$ and ${\alpha}_{\mathrm{1400MHz}}^{\mathrm{144MHz}}=-0.7$ used (Gasperin et al., 2018);
  • detected in FIRST and LoTSS: $S_{\mathrm{obs}}=S_{\mathrm{FIRST}}$ and $\mathtt{alpha\_144\_1400}$ used;
  • not detected in FIRST and LoTSS: the total flux density of VLASS is scaled to that at 1.4 GHz assuming ${\alpha}_{\mathrm{3GHz}}^{\mathrm{1.4GHz}}=-0.7$, and then ${L}_{\mathrm{1.4GHz}}$ is calculated further assuming ${\alpha}_{\mathrm{1400MHz}}^{\mathrm{144MHz}}=-0.7$.
The spectral luminosities will be discussed in detail in § 4.3. The calculations and discussions of radio AB magnitudes ($\mathtt{mag\_\{survey\}}$) and radio loudness ($\mathtt{r\_obs}$) will be presented in § 4.1 and § 4.3, respectively. A full list and details of the columns are given in Table LABEL:tab:_catalog_columns.
Figure 9: Left: VLASS 3 GHz source counts as a function of the AB radio magnitude t. Source counts at $15.511.5<t<$ 15.5 are fitted with a linear function, and the best fits are indicated by the colored line, while the solid ones are fittings in this work and dashed ones are from the literature. The black circles represent UNIONS-VLASS sources, black squares represent the “all” VLASS Epoch 2 catalog with criteria (except $\mathtt{NN\_dist}>30\arcsec$) in § 2.2 applied, and golden crosses represent UNIONS-VLASS sources detected in FIRST, and the size of each sample is shown in the parenthesis. Poisson errors in the sampling bins are represented by vertical bars. Right: Radio number counts normalized to the static Euclidean geometry.
4 Radio and Optical Number Counts
We here summarize the number counts and basic radio properties of AGNs in the UNIONS-VLASS catalog, with the comparisons from previous radio galaxy catalogs in the literature.
4.1 UNIONS-VLASS 3 GHz Radio Source Counts
Radio source number counts are often normalized to Euclidean geometry and examined on the plane of flux density ($S$) and $S^{2.5}\ dN/dS$, where $N$ is the number of sources within each flux density bin (right panel of Figure 9). In this work, we use the plane of the radio AB magnitude t and the number counts N in each magnitude bin to make direct comparisons with other catalogs at 1.4 and 3 GHz. Here, the radio AB magnitude is defined as:
$t\ [\mathrm{AB\ mag}]=-2.5\log\left({{F}_{\mathrm{int}}/3631\ \mathrm{Jy}}\right).$ (4)
The $3\sigma\ (=0.36$ mJy) detection limit of VLASS then corresponds to 17.5 and the $5\sigma\ (=1$ mJy) detection limit of FIRST then corresponds to 16.4.
In the left panel of Figure 9, we show radio source counts for 146,212 “clean” UNIONS-VLASS radio galaxies (black circles). We make a UNIONS-VLASS subsample for 79,638 sources that are also detected in FIRST (golden crosses) and compare it with 1.4 GHz radio galaxy catalogs employing FIRST (Ivezić et al., 2002; Yamashita et al., 2018). We also apply the criteria described in § 2.2, except $\mathtt{NN\_dist>30\arcsec}$, to the entire VLASS2 catalog, resulting in 1,873,519 sources (black squares). We compare this “all” VLASS2 sample with the “clean” VLASS2 sample that contains 1,380,707 sources used for the cross-match. The distributions of these samples at $15.511.5<t<$ 15.5 are fitted with a linear function and are indicated by the solid lines. The best fits of the data collected from the literature are indicated by the dashed lines and are summarized in Table 2.
The trends between “all” VLASS (blue solid line) and UNIONS-VLASS sources (black solid line) are in rough consistency. The global difference and the larger discrepancy at the bright end are attributed to the fact that we have applied $\mathtt{NN\_dist>30\arcsec}$ to VLASS2 in the pre-cross-match stage. This criterion may remove a significant fraction of low-$z$ sources that have bright radio lobes and hotspots identified as adjacent components, contributing to the number counts at the bright end. If we compare UNIONS-VLASS with the “clean” VLASS2 used for the cross-match, then this discrepancy diminishes, and the global difference is consistent with the matching rate of $0.594$.
Compared with 1.4 GHz HSC- and SDSS-FIRST source counts (golden and cyan dashed lines), our 3 GHz UNIONS-VLASS sources (black circles) have a slightly steeper increase at $15.511.5<t<$ 15.5. This could be explained by our choice of $\mathtt{NN\_dist>30\arcsec}$. The best fit of the subsample of our UNIONS-VLASS sources with FIRST detections (golden solid line) has a slope in general consistency with previous studies. This comes from the fact that the sky coverage of UNIONS-VLASS is beyond the survey area of FIRST, thus resulting in an underestimate of the real number counts for our UNIONS-VLASS-FIRST sample.
We further compare the number counts between VLASS2 and FIRST (dashed violet line). Since VLASS2 and FIRST are operated at different frequencies, we scale the 3 GHz flux densities of the “all” VLASS2 adopting $\alpha\mathrm{{}^{FIRST}_{VLASS}}=-0.7$, and the best fit of VLASS2 scaled is indicated by the green solid line. The global amplitude of the scaled “all” VLASS is elevated, whereas the slope becomes flatter. On the other hand, although the “clean” VLASS2 is in general sublinear to FIRST, these two surveys share a consistent trend. These results clearly reflect the complexity of spectral indices of real radio sources that they cannot always be described by a simple power-law distribution (see § 5.1 for details).
Figure 10: Optical UNIONS $r-$band number counts of 146,212 UNIONS-VLASS radio galaxies with valid $r-$band photometry. The blue dots show distributions of sources in each of the randomly selected 2,000 UNIONS tiles, the magenta circles indicate the median of 2,000 UNIONS tiles at each photometry bin, and the black circles represent UNIONS-VLASS sources. The Poisson errors in the sample are represented by vertical bars. The gold solid line indicates the best fit for UNIONS-VLASS sources.
Table 2: Fitting Results of Radio and Optical Number Counts
$a$ $b$
All VLASS $-3.509\ (\pm 0.269)$ $0.317\ (\pm 0.018)$
UNIONS-VLASS $-4.870\ (\pm 0.067)$ $0.391\ (\pm 0.005)$
UNIONS-VLASS-FIRST $-4.779\ (\pm 0.142)$ $0.371\ (\pm 0.010)$
$r-$band $-0.932\ (\pm 0.067)$ $0.074\ (\pm 0.003)$
The best fits of radio source counts in Figure 9 and optical r-band source counts in Figure 10. The fitting function is $\log(N/[\mathrm{mag^{-1}\ deg^{-2}}])=a+bt\ \mathrm{[AB\ mag]}$. The numbers in the parentheses represent $1\sigma$ uncertainties in the fitting results.
4.2 UNIONS-VLASS Optical Source Counts
We investigate the differential counts in $r-$band for our 146,212 “clean” UNIONS-VLASS sources, as shown in Figure 10 by black circles. We also plot the optical counts of UNIONS by randomly selecting 2000 tiles in the NGP. The fall-offs at the faint end (${r}_{\mathrm{ABmag}}\geq 24$) of UNIONS-VLASS and UNIONS are highly consistent, suggesting that the limitation in associating radio galaxies with faint host galaxies is the depth of the optical survey. At the bright end (${r}_{\mathrm{ABmag}}\lesssim 17$), the optical number counts of UNIONS-VLASS sources decrease with the brighter $r-$band magnitude, whereas such a fall-off is not observed in the UNIONS. Additionally, despite the rarity of RL quasars, consistent trends are found between HSC data and HSC-FIRST RGs (Yamashita et al., 2018). Therefore, this fall-off in UNIONS-VLASS is unlikely due to saturation in the $r-$band and the lack of quasars, but arises from an elimination of the brightest nearby radio sources in “clean” VLASS2. For UNIONS-VLASS sources with $19\leq{r}_{\mathrm{ABmag}}\leq 24$, the best fit is indicated by the golden solid line. The source count has a weak dependency on the $r-$band magnitude throughout this range, despite an increasing completeness of UNIONS along with the decreasing $r-$band brightness. This implies a decreasing detectability of radio galaxies in the UNIONS catalog with an increasing apparent optical magnitude: from $0.26\%$ at $r=19$ to $0.01\%$ at $r=24$. Such a trend is also confirmed by VLASS RGs in Euclid Deep Fields-North and Fornax (reaching ${f}_{\mathrm{matches}}\sim 90\%$) down to a VIS magnitude of 28 (Zhong et al. in prep.). This could be the joint consequence of: (1) the Malmquist effect (Miley & De Breuck, 2008) subject to the detection limit of VLASS as the redshift increases; (2) the intrinsic rarity of radio AGNs at high-$z$, partially driven by the increasing energy density of the cosmic microwave background that enhances inverse Compton scattering of synchrotron-radiating electrons, which further exacerbates the selection bias (Condon & Ransom, 2016; Saxena et al., 2017).
Figure 11: Left: observed radio loudness as a function of the redshift. Right: observed radio loudness as a function of the $r-$band magnitudes. These distributions include 63,019 UNIONS-VLASS radio galaxies with valid $r-$band photometric measurements. The observed radio loudness $\mathcal{R}_{\mathrm{obs}}$ is defined as the ratio of the flux density at 1.4 GHz to that of the $r-$band. The RLAGNs are characterized by $\mathcal{R}_{\mathrm{obs}}>1$ and are indicated by orange crosses, while the dark violet pluses indicate RQ AGNs. The contours in the left panel are only for RL populations. The black dots represent 3,579 radio galaxies from the HSC-FIRST catalog (Yamashita et al., 2018). We also show the detection limits of FIRST and VLASS by orange and blue solid lines, respectively.
Figure 12: Left: 1.4 GHz spectral luminosity as a function of the observed radio loudness. Right: 1.4 GHz spectral luminosity as a function of the $r-$band magnitudes. These distributions include 63,019 UNIONS-VLASS radio galaxies with valid $r-$band photometric measurements. The horizontal black dashed line indicates the RL/RQAGN demarcation of ${L}_{\mathrm{1.4GHz}}\sim 10^{24}\mathrm{\ W\ Hz^{-1}}$ (Ivezić et al., 2002). The colored symbols are the same as Figure 11.
Figure 13: Left: The 1.4 GHz spectral luminosity as a function of photo-z for ${\color[rgb]{0,0,0}63,019}$ UNIONS-VLASS radio galaxies (orange crosses), in comparison with HSC-FIRST AGNs (black dots) and local RLAGNs (cyan triangles). VLASS flux densities at 3 GHz are converted to those at 1.4 GHz, adopting ${\alpha}_{\mathrm{VLASS}}^{\mathrm{FIRST}}$ for those detected in FIRST and adopting $\alpha=-0.7$ for those without FIRST detection. Right: The 3 GHz spectral luminosity as a function of photo-z for UNIONS-VLASS sample, in comparison with VLA 3 GHz survey. We separate ${\color[rgb]{0,0,0}63,019}$ UNIONS-VLASS radio galaxies with valid ${z}_{\mathrm{photo}}$ into two subsamples. ${L}_{\mathrm{3GHz}}$ of those detected in both VLASS and FIRST (${\color[rgb]{0,0,0}41,803}$ in total) are calculated based on $\alpha\mathrm{{}_{VLASS}^{FIRST}}$, and of those without FIRST detection (${\color[rgb]{0,0,0}21,216}$ in total) are calculated using $\alpha=-0.7$. These sources are further divided into RL and RQ populations following the division line ${L}_{\mathrm{1.4GHz}}\sim 10^{24}\mathrm{\ W\ Hz^{-1}}$ (Ivezić et al., 2002), which is indicated by the blue solid line. In both panels, the violet line indicates the $5\sigma\ (=1.0)$ mJy detection limit of FIRST and the black one indicates the $3\sigma\ (=0.36)$ mJy detection limit of VLASS.
4.3 Radio-Loud AGNs
4.3.1 Radio Loudness
To separate RLAGNs from radio-quiet (RQ) populations, the observed radio loudness $\mathcal{R}_{\mathrm{obs}}$ is often used. A widely adopted quantification of $\mathcal{R}_{\mathrm{obs}}$ is defined as the radio-to-optical flux ratio where no $k-$corrections are applied:
$\log\mathcal{R}_{\mathrm{obs}}=\log\left({{F}_{\mathrm{radio}}/{F}_{\mathrm{optical}}}\right)=0.4(m-t),$ (5)
where $m$ is one of the optical magnitudes and $t$ is the AB radio magnitudes in Eq. (3). Although the i-band photometry is commonly chosen for ${F}_{\mathrm{optical}}$, the fact that UNIONS5000 is built upon the $r-$band detections ensures a hundred percent completeness of $\mathcal{R}_{\mathrm{obs}}$ for our catalog. We thus use the $r-$band photometry instead. The median difference between $r-$ and $i-$band photometry, as shown in the left panel of Figure 22, is 0.44 in both UNIONS5000 and UNIONS2000 data releases. Therefore, statistically, the choice of ${r}_{\mathrm{ABmag}}$ has insignificant impacts on the final classifications. For ${F}_{\mathrm{radio}}$, we use the 1.4 GHz flux density. The 1.4 GHz flux densities of UNIONS-VLASS sources with FIRST detection are calculated using FIRST flux densities. Those not detected in FIRST have their VLASS 3 GHz flux densities scaled to those at 1.4 GHz using ${\alpha}_{\mathrm{3GHz}}^{\mathrm{1.4GHz}}=-0.7$. RLAGNs are then selected with $\log\mathcal{R}_{\mathrm{obs}}>1$. In this manner, for 146,212 sources in the “clean” UNIONS-VLASS catalog, 138,266 ($\sim 94.6\%$) are RL and 7,946 ($\sim 5.4\%$) are RQ populations.
4.3.2 UNIONS-VLASS Subsample with Valid Photo-z
To better understand RLAGNs across cosmic time, we focus on the subsample of ${\color[rgb]{0,0,0}63,019}$ radio galaxies with valid UNIONS photo-$z$. In this subsample, 58,661 are RL and 4,358 are RQ populations based on the $\mathcal{R}_{\mathrm{obs}}$ criterion. We show the distribution of $\mathcal{R}_{\mathrm{obs}}$ as a function of the photometric redshift in the left panel of Figure 11. Compared to the HSC-FIRST catalog that has $\mathcal{R}_{\mathrm{obs}}\sim 4$ at ${z}_{\mathrm{photo}}\sim 2-4$ (Yamashita et al., 2018), the bulk of our UNIONS-VLASS radio galaxies at this redshift range have $\mathcal{R}_{\mathrm{obs}}\sim 2-3$, and there is no evolution of $\mathcal{R}_{\mathrm{obs}}$ within this redshift bin. This may be explained by the fact that the HSC-FIRST catalog uses ${i}_{\mathrm{ABmag}}$ rather than ${r}_{\mathrm{ABmag}}$, and because of the greater depth of the HSC $i-$band, Yamashita et al. (2018) captured more optically faint sources. In the panel showing $\mathcal{R}_{\mathrm{obs}}-{r}_{\mathrm{ABmag}}$ (right panel of Figure 11), we also plot the detection limit of VLASS and FIRST. We see that at ${i}_{\mathrm{ABmag}}\gtrsim 19.5$, all radio galaxies are radio-loud, and the RQ populations only exist at ${r}_{\mathrm{ABmag}}$ brighter than this magnitude, which clearly reflects the selection bias of RLAGNs based on $\mathcal{R}_{\mathrm{obs}}$, that is, this criterion favors optically faint galaxies.
To investigate whether the current abundance of RLAGNs in UNIONS-VLASS is due to the faintness of the optical counterparts, we adopt an alternative approach to selecting RLAGNs via the spectral luminosity threshold ${L}_{\mathrm{radio}}\sim 10^{24}\mathrm{\ W\ Hz^{-1}}$ established from FIRST RL/RQAGNs (Ivezić et al., 2002). This threshold depends on the reduced Hubble constant and we adopted ${h}_{\mathrm{50}}$ for consistency between works, and ${L}_{\mathrm{radio}}\sim 2\times 10^{24}\mathrm{\ W\ Hz^{-1}}$ if ${h}_{\mathrm{70}}$ is used instead, leading to no significant difference. For ${\color[rgb]{0,0,0}63,019}$ UNIONS-VLASS radio galaxies with valid ${z}_{\mathrm{photo}}$, ${\color[rgb]{0,0,0}49,231}$ are RLAGNs with ${L}_{\mathrm{radio}}>10^{24}\mathrm{\ W\ Hz^{-1}}$ There are ${\color[rgb]{0,0,0}49,012}$ sources commonly identified as RLAGNs via selections based on ${L}_{\mathrm{1.4GHz}}$ and $\mathcal{R}_{\mathrm{obs}}$. Fractionally, the consistency is $\sim 99.6\%$ (49,012/49,231) for ${L}_{\mathrm{1.4GHz}}$ and $\sim 83.6\%$ (49,012/58,661) for $\mathcal{R}_{\mathrm{obs}}$, respectively.
We show ${L}_{\mathrm{1.4GHz}}$ as a function of $\mathcal{R}_{\mathrm{obs}}$ and ${r}_{\mathrm{ABmag}}$ in Figure 12. It is revealed that RQAGNs classified from $\mathcal{R}_{\mathrm{obs}}<1$ generally are weak in ${L}_{\mathrm{1.4GHz}}$ in comparison with RLAGNs. Nonetheless, some RQAGNs are brighter than ${L}_{\mathrm{1.4GHz}}=10^{25}\mathrm{\ W\ Hz^{-1}}$ with ${i}_{\mathrm{ABmag}}=16-18$, suggesting that they could be RL quasars. There is a clear correlation between ${L}_{\mathrm{1.4GHz}}\ (\sim 10^{23}-10^{24}\mathrm{\ W\ Hz^{-1}})$ and $\mathcal{R}_{\mathrm{obs}}\ (\sim 0-2)$. Similarly, a weak correlation also exists for ${L}_{\mathrm{1.4GHz}}\ (\sim 10^{23}-10^{24}\mathrm{\ W\ Hz^{-1}})$ and ${r}_{\mathrm{ABmag}}\ (\sim 16-18)$. These correlations are likely associated with the cosmic evolution of $\mathcal{R}_{\mathrm{obs}}$ at $z<1$ (left panels of Figure 11 and Figure 21). On the other hand, beyond the RL-RQ demarcation of ${L}_{\mathrm{radio}}\sim 10^{24}\mathrm{\ W\ Hz^{-1}}$, the spectral luminosity has no evolution along ${r}_{\mathrm{ABmag}}\sim 22-26$.
We further plot 1.4 GHz and 3 GHz spectral luminosities in the left and right panels of Figure 13, respectively. For those detected in FIRST, ${L}_{\mathrm{3GHz}}$ is calculated adopting ${\alpha}_{\mathrm{VLASS}}^{\mathrm{FIRST}}$, while those without detection adopt a spectral index of $-0.7$. We overlay the RL and RQ demarcation on the distribution of ${L}_{\mathrm{1.4GHz}}$ and scale it to that of ${L}_{\mathrm{3GHz}}$ adopting ${\alpha}_{\mathrm{3GHz}}^{\mathrm{1.4GHz}}=-0.7$. Limited by the sensitivity, both VLASS and FIRST cannot reach the depth required to detect RQAGNs at $z>1$. Therefore, our UNIONS-VLASS radio galaxies $z>1$ are (almost) radio-loud, no matter whether they are identified based on $\mathcal{R}_{\mathrm{obs}}>1$ or ${L}_{\mathrm{1.4GHz}}>10^{24}\mathrm{\ W\ Hz^{-1}}$.
Figure 14: The radio color-color diagram or spectral index map of ${\alpha}_{\mathrm{FIRST=1400MHz}}^{\mathrm{LoTSS=144Mhz}}$ and ${\alpha}_{\mathrm{VLASS=3000MHz}}^{\mathrm{FIRST=1400Mhz}}$ for 30,091 UNIONS-VLASS sources with FIRST and LoTSS detections. The darker and larger marker represents a higher redshift. The spectral shapes are shown in the four quadrants, where Q1 represents upturned, Q2 represents inverted, Q3 represents power-law, Q4 represents peaked spectrum, respectively (see texts for details).
5 Discussions
5.1 Radio Color-color Diagram
The spectral shape of the synchrotron radiation and the projected linear sizes of the synchrotron-radiating regions offer implications for the evolution stage and environment of an RLAGN (Miley & De Breuck, 2008; Murgia et al., 2011; Gasperin et al., 2018; Hardcastle & Croston, 2020; Zhong et al., 2023). To characterize the spectral shapes of our UNIONS-VLASS sources, we show the radio color-color diagram (or spectral index map) between ${\alpha}_{\mathrm{FIRST=1400MHz}}^{\mathrm{LoTSS=144Mhz}}$ and ${\alpha}_{\mathrm{VLASS=3000MHz}}^{\mathrm{FIRST=1400Mhz}}$ in Figure 14, which includes 30,091 sources detected in VLASS, FIRST, and LoTSS with $\mathtt{alpha\_\{frequency1\}\_\{frequency2\}\_err}<0.2$. The expected spectral shapes covering 144 MHz to 3 GHz are indicated by the orange lines and can be divided into four quadrants (labeled Q14141-41 - 4).
Sources located in quadrant 1 have an upturned spectrum, which could be explained by a burst of synchrotron radiation subsequent to the FIRST observations conducted mostly in the 1990s, or the inhomogeneity of the radio regions observed (Condon & Ransom, 2016). Sources located in quadrant 2 have an inverted spectrum. One possible explanation is the sudden increase in the radio brightness. More commonly, it is due to the fact that synchrotron self-absorption (SSA; e.g., Tingay & de Kool, 2003; Nyland et al., 2020) dominates synchrotron radiation. Sources located in quadrant 3 have the most commonly observed power-law distribution of the non-thermal bremsstrahlung. These sources may have a steepening of the spectral index, forming the curved power-law spectrum as a result of synchrotron losses and/or inverse-Compton losses (Dermer & Menon, 2009; Blandford et al., 2019). Sources located in quadrant 4 have a peaked frequency (${\nu}_{\mathrm{p}}$) inasmuch an equilibrium between the brightness and kinetic temperatures of the synchrotron-radiating electrons (Longair, 2011). The corresponding radio galaxies could be GHz-peaked ($0.5\leq{\nu}_{\mathrm{p}}\leq 5$ GHz) or high-frequency-peaked (${\nu}_{\mathrm{p}}>5$ GHz) sources depending on their redshifts, which comprises the so-called Peaked-Spectrum (PS) sources that are young radio galaxies (see O’Dea & Saikia 2021 and references therein). Sources with inverted spectrum between 144 MHz and 3 GHz (in quadrant 2) may also have peaked/turnover frequencies above 3 GHz, making them PS sources.
Figure 15: Spectral index between LoTSS 144 MHz and VLASS 3 GHz as function of the UNIONS BPZ photo-$z$ and the projected linear size considering the deconvolution of the beam size based on VLASS observations for 16,078 UNIONS-VLASS sources reliable in ${z}_{\mathrm{photo}}$ and detected in LoTSS. The orange line indicates ${\alpha}_{\mathrm{3000MHz}}^{\mathrm{144MHz}}=0.5$. The black line indicates ${\alpha}_{\mathrm{3000MHz}}^{\mathrm{144MHz}}=-1.3$, below which the sources are classified as USS sources. The blue line indicates ${\alpha}_{\mathrm{3000MHz}}^{\mathrm{144MHz}}=-0.5$ and projected linear size smaller than 20 kpc, and the enclosed region defines CSS sources.
Figure 16: Spectral index between FIRST 1.4 GHz and VLASS 3 GHz as function of the UNIONS BPZ photo-$z$ and the projected linear size considering the deconvolution of the beam size based on VLASS observations for ${\color[rgb]{0,0,0}15,906}$ UNIONS-VLASS sources detected in FIRST and are valid in ${z}_{\mathrm{photo}}$. The colored lines are the same as Figure 15, and an additional violet line indicates the SSA-dominated inverted spectral index of $+2.5$.
The sizes of the synchrotron-emitting volume matter when it comes to the classification of radio galaxies. We further show the spectral index ${\alpha}_{\mathrm{VLASS=3000MHz}}^{\mathrm{LoTSS=144MHz}}$ (${\color[rgb]{0,0,0}16,078}$ UNIONS-VLASS-LoTSS sources) and ${\alpha}_{\mathrm{VLASS=3000MHz}}^{\mathrm{FIRST=1400MHz}}$ (${\color[rgb]{0,0,0}15,906}$ UNIONS-VLASS-FIRST sources) as a function of photo-$z$ and the VLASS projected linear size of the radio regions deconvolved from the beam size in Figure 15 and 16, respectively, considering only those with reliable spectral index estimates (uncertainties smaller than 0.2). The black solid line indicates a spectral index of $-1.3$, and sources with a spectral slope steeper than this value are called Ultra-Steep-Spectrum (USS; De Breuck et al. 2000). This could be due to a significant synchrotron ageing, suggestive of well-evolved RLAGNs (van Breugel et al., 1984; Turner, 2018). USS is often used to select high-$z$ RLAGNs as energy losses could be dominated by inverse-Compton scattering, as well as being attributed to the redshifting of the frequency (Carilli & Yun, 1999; Morabito & Harwood, 2018). The blue solid line indicates a spectral index of $-0.5$. If a source has $\alpha<-0.5$ at GHz regime and its projected linear size is more compact than 20 kpc, it is then classified as Compact-Steep-Spectrum (CSS; Callingham et al. 2017; Nyland et al. 2020; Orienti et al. 2023). CSS sources are often thought to be RLAGNs at an early evolution stage, otherwise, they are actually evolved populations while their jets are confined by the ISM of their host galaxies (Murgia, 2003; Mukherjee et al., 2016). Some sources lie above ${\alpha}_{\mathrm{3GHz}}^{\mathrm{1.4GHz}}=2.5$, which is beyond the inverted spectral slope expected to be under the dominance of SSA. They are likely to be transients that undergo a sudden outburst of radio jets.
In this work, we only briefly report the spectral index distribution of our UNIONS-VLASS sources. A detailed and thorough investigation of the radio classifications of our sources will be presented in the future work (Zhong et al. in preparation).
Figure 17: Color-color diagrams to select dropout sources. The upper and lower panels show selections for g- ($53<z<$ 5) and r-dropout ($64<z<$ 6) sources, respectively. The black lines represent the criteria adopted to select dropouts (Eqs. (5)-(6)). The blue dots are all UNIONS-VLASS sources with valid photometry in the corresponding two-color diagram. The contours are in $1\sigma$, $2\sigma$, and $3\sigma$ confidence levels.
5.2 High-$z$ Candidates with $z\geq 3$
As introduced in § 5.1, the USS characterized by $\alpha<-1.3$ is an efficient technique to select HzRGs. However, in addition to these USS sources, many HzRGs have been discovered with relatively flat spectral index (e.g., ${\alpha}_{\mathrm{1.4GHz}}^{\mathrm{325MHz}}=-0.75$ for J163912.11+405236.5 at $z=4.88$, Jarvis et al. 2009; ${\alpha}_{\mathrm{1.4GHz}}^{\mathrm{325MHz}}=-1.1$ for HSC J083913.17+011308.1 at $z=4.72$, Yamashita et al. 2020; see also Miley & De Breuck 2008). As a consequence, a sole reliance on USS could lead to a bias in the statistical physical properties of the HzRGs since the diverse nature could be over-represented by USS.
An alternative approach to select HzRGs, which can circumvent the selection bias arising from the uncertain real spectral index distributions, is the Lyman break technique commonly adopted in the search for high-$z$ galaxies. The effectiveness of this technique in discovering HzRGs has been proven by Capetti et al. (2025), who identified 39 $g-$dropout HzRGs expected to be at $4.53<z<$ 4.5 over an area of $\sim 2,000\mathrm{\ deg^{2}}$, and by Yamamoto et al. (2025), who constructed a photometric sample of 146 HzRGs at $z\sim 4$ within a total survey area of $\sim 560\mathrm{\ deg^{2}}$.
The Lyman break color selection technique selects sources that exhibit a clear Lyman break and blue UV continuum in their optical SEDs covered by broadband (ugriz) photometry. Following Hildebrandt et al. (2009) that have used a similar filter set, g-dropouts ($53<z<$ 5) are selected according to
$g-r>1.0,$ (6)
$r-i<1.0,$
$g-r>1.5(r-i)+0.8,$
and r-dropouts ($64<z<$ 6) are extracted adopting
$r-i>1.2,$ (7)
$i-z<0.7,$
$r-i>1.5(i-z)+1.0.$
In each dropout selection, three adjacent bands have to be used. We therefore only select dropout candidates if the photometric measurements of these bands are not flagged and have valid AB magnitudes (see § 3.3). Additionally, for g-dropouts, the observation can barely detect fluxes prior to the Lyman break, thus the u-band should be fainter than the detection limit with $\mathtt{MAG\_GAAP\_u==99}$, that is, there is no successful photometric measurement of the corresponding source. Likely, this criterion of u-band for g-dropouts should be applied to both u- and g-bands for r-dropouts. Moreover, to ensure these high-$z$ candidates are real detections, we further require their ${i}_{\mathrm{ABmag}}$ to be brighter than the $10\sigma$ detection limit 23.8. Furthermore, to avoid foreground contaminants, we apply a magnitude cut of ${i}_{\mathrm{ABmag}}\geq 23$.
We show g- and r-dropout candidates in the upper and lower panels of Figure 17, respectively. The above criteria produce 4,049 valid sources for the r-dropouts, and only 12 of them can be at $z\sim 5$, which is reasonable considering the limited depth of UNIONS. Of ${\color[rgb]{0,0,0}63,019}$ UNIONS-VLASS sources with reliable BPZ photo-$z$, there are only 2 at $z>4.5$, comparable with the r-dropout selections. The number of g-dropout-selected sources is $477$, which exceeds the number counts of 53 expected from reliable $4.53.5<{z}_{\mathrm{photo}}<$ 4.5 estimates by almost one order of magnitude.
This potential excess in the number counts of g-dropouts – if they are not bona fide sources at $z\sim 4$ – could be explained by the contamination of foreground sources such as red elliptical galaxies at intermediate redshifts, whose ${D}_{\mathrm{4000}}$ break due to the absorption of old stellar populations mimic the Lyman break feature. The photometric errors of these foreground contaminations satisfy the selection criteria of the g-dropout (e.g., Ono et al., 2018). For Subaru HSC, which has comparable optical number counts and similar efficient broadband wavelengths to those of UNIONS, Ono et al. (2018) found the contamination rate in the $i-$band magnitude bin of 22.5 can reach $\sim 68\%$ and decreases to $\sim 41\%$ at ${i}_{\mathrm{ABmag}}=23.1$ and to $\sim 35\%$ at ${i}_{\mathrm{ABmag}}=23.7$ for g-dropouts. Therefore, even if we only select UNIONS galaxies at the faint end of $i-$band (${i}_{\mathrm{AB}}>23$), the contamination problem can still persist. Similarly, for $r-$dropouts, Ono et al. (2018) reported contaminate rates of approximately $23\%$ at ${z}_{\mathrm{ABmag}}=22.3$ and around $59\%$ at ${z}_{\mathrm{ABmag}}=23.5$, implying that merely $<5$ $r-$dropouts could be true positives if applying these estimates to UNIONS-VLASS. Nonetheless, under the optimistic assumption of HSC-level contamination rates, our UNIONS-VLASS radio galaxy catalog is expected to yield more than 200 HzRGs at $z\sim 4$, offering strong potential to significantly enhance the statistical robustness of studies on HzRGs.
All in all, these expectations for high-z RLAGNs are subject to the reliability of the photometry of UNIONS. A deeper investigation into the HzRGs will be conducted in future work using spectral energy distribution fittings combining IR data (Zhong et al. in preparation).
5.3 Caveat to the Catalog
The effect of applying $\mathtt{NN\_dist>30^{\prime\prime}}$ to VLASS2 is clearly reflected in the projected linear size, that is, the majority of our RGs are compact (${l}_{\mathrm{s}}\leq 20$ kpc), which means our UNIONS-VLASS RGs have the tendency to be young and evolving populations. This also leads to a potential bias of our HzRG populations. As they are often selected by the USS feature in previous studies, there are only a few of them showing USS at $z>2$ amongst $15,906$ UNIONS-VLASS-FIRST sources with valid ${\alpha}_{\mathrm{VLASS}}^{\mathrm{FIRST}}$. This global trend of being compact for our RGs is more a result of the limited search radius of $1.8^{\prime\prime}$, As shown in Figure 2, in $\mathtt{NN\_dist<30^{\prime\prime}}$, there is no concentration of radio sources with small offsets ($\leq 1^{\prime\prime}$) from their optical counterparts. To maintain the current cleanliness of UNIONS-VLASS catalog, even if $\mathtt{NN\_dist<30^{\prime\prime}}$ is not adopted, we do not expect a significant increase in the number counts of bright and extended radio sources. We, thereby, remind the readers to be cautious about the intrinsic and unresolved bias arising from the design of this work. Nonetheless, the vast size of this catalog still ensures statistical studies of radio AGNs across various evolution stages, offering valuable insights into the connections between radio AGNs and their host galaxies.
6 Summary
We presented the results of searching for the optical counterparts of radio sources using the grand UNIONS catalog and the VLASS catalog. This marks the first step of the ongoing project, UNIONS-VLASS radio galaxies in the Euclid wide survey area (UNVEIL). The current angular separation-based cross-match between UNIONS and VLASS found 146,212 radio galaxies/quasars spanning a wide area of $\sim 4,200\ \mathrm{deg^{2}}$ in the northern and southern Galactic caps. The fraction of matches reaches 0.594, beyond the previous FIRST 1.4 GHz radio galaxy catalog based on the SDSS and Subaru HSC catalogs.
Setting UNIONS-VLASS as the reference catalog, we further performed cross-match with FIRST 1.4 GHz and LoTSS 144 MHz catalogs, yielding 79,638 UNIONS-VLASS-FIRST, 101,671 UNIONS-VLASS-LoTSS, and 64,672 UNIONS-VLASS-FIRST-LoTSS radio galaxies, respectively. We identified radio-loud AGNs based on the observed radio loudness ($\mathcal{R}_{\mathrm{obs}}$) defined as the ratio of 1.4 GHz radio to r-band optical flux densities. Of 146,212 “clean” UNIONS-VLASS radio galaxies, 138,266 ($\sim 94.6\%$) are RL and 7,496 ($\sim 5.4\%$) are RQ populations. Alternatively, RLAGNs can be selected by their 1.4 spectral luminosities ${L}_{\mathrm{1.4GHz}}>10^{24}\mathrm{\ W\ Hz^{-1}}$. Under this condition, at $z\gtrsim 1$, both VLASS and FIRST cannot detect any RQ populations. RLAGNs selections based on $\mathcal{R}_{\mathrm{obs}}$ and ${L}_{\mathrm{1.4GHz}}$ are common for 49,012 out of 63,019 sources with valid UNIONS photometric redshift estimates.
We presented radio color-color (${\alpha}_{\mathrm{1400MHz}}^{\mathrm{144MHz}}-{\alpha}_{\mathrm{3000MHz}}^{\mathrm{1400MHz}}$) based on LoTSS, FIRST, and VLASS observations, and further plotted spectral indices as a function of redshift and the projected linear size of the synchrotron-emitting volume. We found a great diversity of radio classifications, including Compact-Steep-Spectrum and Peaked-Spectrum sources that are likely young radio galaxies, Ultra-Steep-Spectrum sources which may have significantly evolved, and inverted and upturned spectrum that may be linked to transient sources. There are $\sim 10,000$ additional radio galaxies corresponding to more than one optical counterpart. These sources are merger candidates that host radio galaxies. The radio classifications/RLAGN evolution and merger-RLAGN connections will be further studied in future works.
The UNIONS internal data release used for this work does not cover the full UNIONS footprint in both NGC and SGC. In addition, UNIONS is an ongoing survey that will eventually cover the entire northern hemisphere. Moreover, the VLASS Epoch 3 observation has just been completed, and the concatenated catalog reaching a great depth has not been released. Therefore, our UNIONS-VLASS catalog has great potential for expansion. The full UNIONS footprint will be covered by the Euclid wide survey, and this work serves as a valuable pavement for the future Euclid radio galaxy catalog. Our UNIONS-VLASS radio galaxy catalog will be observed in the Subaru PFS filler program. By virtue of the broad spectroscopic coverage ($\sim 3000-20000\AA$) and imaging data, the abundance of RLAGNs in the UNIONS-VLASS catalog will profoundly facilitate multifaceted statistical studies of radio galaxies and beyond.
Acknowledgments
We wish to thank Takuji Yamashita and Hongming Tang for the discussions. YZ is supported by Japan Society for the Promotion of Science Research Fellowship for Young Scientists. This work is supported by Japan Society for the Promotion of Science KAKENHI (23H00131; Akio K. Inoue, 25K01043; K. Ichikawa, 24K22894; M. Onoue).
We are honored and grateful for the opportunity of observing the Universe from Maunakea and Haleakala, which both have cultural, historical and natural significance in Hawaii. This work is based on data obtained as part of the Canada-France Imaging Survey, a CFHT large program of the National Research Council of Canada and the French Centre National de la Recherche Scientifique. Based on observations obtained with MegaPrime/MegaCam, a joint project of CFHT and CEA Saclay, at the Canada-France-Hawaii Telescope (CFHT) which is operated by the National Research Council (NRC) of Canada, the Institut National des Science de l’Univers (INSU) of the Centre National de la Recherche Scientifique (CNRS) of France, and the University of Hawaii. This research used the facilities of the Canadian Astronomy Data Centre operated by the National Research Council of Canada with the support of the Canadian Space Agency. This research is based in part on data collected at Subaru Telescope, which is operated by the National Astronomical Observatory of Japan. Pan-STARRS is a project of the Institute for Astronomy of the University of Hawaii, and is supported by the NASA SSO Near Earth Observation Program under grants 80NSSC18K0971, NNX14AM74G, NNX12AR65G, NNX13AQ47G, NNX08AR22G, 80NSSC21K1572 and by the State of Hawaii.
The National Radio Astronomy Observatory is a facility of the National Science Foundation operated under cooperative agreement by Associated Universities, Inc. CIRADA is a partnership between the University of Toronto, the University of Alberta, the University of Manitoba, the University of British Columbia, McGill University and Queen’s University, in collaboration with National Research Council of Canada, the US National Radio Astronomy Observatory, and Australia’s Commonwealth Scientific and Industrial Research Organization. CIRADA is funded by a grant from the Canadian Foundation for Innovation 2017 Innovation Fund (Project 35999) and the by the Provinces of Ontario, British Columbia, Alberta, Manitoba and Quebec.
This research used the facilities of the Canadian Astronomy Data Centre operated by the National Research Council of Canada with the support of the Canadian Space Agency.
Appendix A Radio Morphology of Sources with $\mathtt{NN\_dist}<30^{\prime\prime}$
We randomly choose 36 VLASS radio sources with $\mathtt{NN\_dist}<30^{\prime\prime}$ and show their Epoch 2 QL images in
Figure 18: VLASS Epoch 2 QL images of VLASS sources with $\mathtt{NN\_dist}<30^{\prime\prime}$. Cyan crosses represent the centroids of VLASS sources and black pluses represent the centroids of UNIONS sources cross-matched within a search radius of $2.5^{\prime\prime}$.
Appendix B Additional Astrometric Comparisons
We show the absolute values of astrometric offsets along R.A. for 146,212 UNIONS-VLASS radio galaxies cross-matched via angular separations $\leq 1.5\arcsec$ between UNIONS and VLASS sources as a function of the declination in the left panel of Figure 19. It makes clear that large astrometric offsets along R.A. are attributed to the projection effect, and this worsens as the declination increases. For previous studies based on VLA FIRST and adopting positional cross-match, this situation is barely considered since the sky coverage of FIRST is limited to $\mathrm{DEC<65^{\circ}}$, and the projection effect does not lead to a significant excess of the astrometric offset along R.A. However, for VLASS that approaches the north pole, an angular separation-based cross-match is preferred over the positional one, otherwise, a small fraction of the radio sources can be missed in the UNIONS-VLASS ($\sim 4\%$), as well as in other future catalogs using VLASS.
Figure 19: Left: the offsets along the R.A. as a function of the declination for 146,212 “clean” UNIONS-VLASS radio galaxies, where the vertical line indicates the angular separation-based cross-match radius of $1.5\arcsec$. The offset along R.A. suffers from a more severe projection effect at higher declinations. Right: the plane of offset along R.A. versus the offset along declination.
Appendix C UNIONS-VLASS r-band Source Counts for RL and RQAGNs
Figure 20: Optical $r-$band source counts for 63,019 UNIONS-VLASS radio galaxies (circles). In both panels, squares represent RL and crosses represent RQ populations. Left: the UNIONS-VLASS sources are divided into RL and RQAGNs based on the radio loudness, where RLAGNs have $\mathcal{R}_{\mathrm{obs}}>1$. Right: RL and RQAGNs are separated via the 1.4 GHz spectral luminosity, where RLAGNs have ${L}_{\mathrm{1.4GHz}}>10^{24}\mathrm{\ W\ Hz^{-1}}$.
We divide 63,019 UNIONS-VLASS sources with valid photo-$z$ into RL (squares) and RQ (crosses) subsample based on the observed radio loudness ($\mathcal{R}_{\mathrm{obs}}$; see § 4.3 for details) and show the optical $r-$band source counts for the two populations, respectively, in the left panel of Figure 20. We see that RLAGNs dominate the AGN populations whose optical counterparts lie above $r\gtrsim 20$. On the other hand, the brightest optical counterparts are almost host galaxies to RQAGNs, and there are no RQAGNs in optically faint galaxies ($r>20$). This is consistent with the distribution of the radio loudness against ${r}_{\mathrm{ABmag}}$. For optically luminous sources, their large optical flux densities make it more difficult for them be radio-loud even though they are powerful in synchrotron radiation. In this manner, RLAGNs are preferentially associated with massive systems with inactive star-forming actives, while RQAGNs are more likely to reside in systems with intense star formation.
We further divide RL and RQ populations based on their 1.4 GHz spectral luminosities (see § 4.3 for details). We see that RQAGNs are still the dominant populations for galaxies at ${r}_{\mathrm{ABmag}}<18.5$. However, at ${r}_{\mathrm{ABmag}}>20$, the source counts of RQAGNs do not decrease as steeply as the trend based on $\mathcal{R}_{\mathrm{obs}}$, which is similar to that of radio quasars (see, e.g, Yamashita et al. 2018). With the increasing $r-$band magnitude, there is a lower possibility of finding RQAGNs compared to RL populations. In summary, from this aspect, both RL and RQAGNs are hosted by a large variety of galaxy populations.
Appendix D Radio-loud AGNs
Star formation can contribute to the non-thermal emission, and rigorously $\log\mathcal{R}_{\mathrm{obs}}>2$ may be adopted to exclude star-forming/starburst galaxies that mimic RLAGNs (e.g., Chiaberge & Marconi, 2011). This reduces our RLAGNs to 109,901 ($\sim 77.0\%$). The spectral luminosity of free-free emission attributed to star formation can be quantified by (Condon, 1992):
$\left({\frac{{L}_{\mathrm{NT}}}{\mathrm{\ W\ Hz^{-1}}}}\right)\sim 5.5\times 10^{20}\left({\frac{\nu}{\mathrm{GHz}}}\right)^{-0.1}\left[\frac{\mathrm{SFR(\mathit{{M}_{\mathrm{star}}}\geq\mathrm{M_{\odot}})}}{\mathrm{M_{\odot}}\ \rm yr^{-1}}\right],$ (D1)
and the non-thermal emission related to SF activities follows
$\left({\frac{{L}_{\mathrm{NT}}}{\mathrm{\ W\ Hz^{-1}}}}\right)\sim 5.3\times 10^{21}\left({\frac{\nu}{\mathrm{GHz}}}\right)^{-0\alpha}\left[\frac{\mathrm{SFR(\mathit{{M}_{\mathrm{star}}}\geq\mathrm{M_{\odot}})}}{\mathrm{M_{\odot}}\ \rm yr^{-1}}\right].$ (D2)
Assuming an upper limit of $\mathrm{SFR}=300\ \mathrm{M_{\odot}}\ \rm yr^{-1}$ for star-forming main sequence galaxies at $z\gtrsim 2$ (Popesso et al., 2023), the corresponding luminosities are ${L}_{\mathrm{T,SF}}\sim 1.5\times 10^{23}\mathrm{\ W\ Hz^{-1}}$ and ${L}_{\mathrm{NT,SF}}\sim 1.3\times 10^{24}\mathrm{\ W\ Hz^{-1}}$. In Figure 13 for ${L}_{\mathrm{1.4GHz}}$ and ${L}_{\mathrm{3GHz}}$ as a function of $z$, this ${L}_{\mathrm{NT,SF}}$ only selects radio galaxies at $z\lesssim 1$ where the host galaxies should have much smaller values of SFR compared to high-z sources even if they are rarely observed starburst systems. The real contributions of star-forming activities to the observed synchrotron radiation should be insignificant ($\lesssim 1.5\times 10^{22}\mathrm{\ W\ Hz^{-1}}$ assuming $\mathrm{SFR}=40\ \mathrm{M_{\odot}}\ \rm yr^{-1}$ at $z\sim 0.5$). Therefore, $\mathcal{R}_{\mathrm{obs}}>1$ can be safely adopted to select RLAGNs for the UNIONS-VLASS sample.
We see correlations between ${L}_{\mathrm{1.4GHz}}-\mathcal{R}_{\mathrm{obs}}$ and $\mathcal{R}_{\mathrm{obs}}-{z}_{\mathrm{photo}}$, and these correlations are likely a propagation of the correlation between ${i}_{\mathrm{ABmag}}-{z}_{\mathrm{photo}}$ shown in the left panel of Figure 21 that reflects the cosmological surface brightness dimming. As a result, high-$z$ sources have higher possibilities of being classified as RLAGNs because of their optical faintness. In the right panel of Figure 21, we plot absolute radio magnitude $M_{t}$ as a function of the observed radio loudness $\mathcal{R}_{\mathrm{obs}}$, where $M_{t}$ is defined as (Ivezić et al., 2002)
$M_{t}=-2.5\log\left({\frac{{L}_{\mathrm{radio}}}{{L}_{\mathrm{AB}}}}\right).$ (D3)
${L}_{\mathrm{radio}}$ is the specific or spectral radio luminosity, and ${L}_{\mathrm{AB}}=4.345\times 10^{13}\mathrm{\ W\ Hz^{-1}}$ is the specific luminosity of a source whose specific flux is 3631 Jy at a distance of 10 pc. Normally, $M_{t}=-25.9$ is chosen as the RL/RQ demarcation, corresponding to ${L}_{\mathrm{1.4GHz}}\sim 10^{24}{h}_{\mathrm{50}}^{\mathrm{-2}}\mathrm{\ W\ Hz^{-1}}$. For 146,212 UNIONS-VLASS sources with valid r-band photometry to calculate $\mathcal{R}_{\mathrm{obs}}$, the two thresholds to select RLAGNs are the same for 83.3%.
Figure 21: Left: UNIONS $r-$band photometry as a function of the photometric redshift. Right: Absolute radio magnitude as a function of the observed radio loudness, where the dashed line indicates the RL/RQ demarcation of $M_{t}=-25.9$. The golden dots represent RLAGNs based on $\mathcal{R}_{\mathrm{obs}}>1$ and the violet ones are RQAGNs defined by $\mathcal{R}_{\mathrm{obs}}<1$.
Appendix E UNIONS2000 versus UNIONS5000 for i-band Photometry
We compare the distributions of the $r-i$ color between UNIONS2000 and UNIONS5000 for 76,454 UNIONS-VLASS radio galaxies produced in both data releases in Figure 22. The median difference between ${r}_{\mathrm{ABmag}}$ and ${i}_{\mathrm{ABmag}}$ is 0.44 in both UNIONS2000 and UNIONS5000. However, in comparison with UNIONS2000, the distribution of $r-i$ in UNIONS5000 is more convergent, which is positively attributed to that the UNIONS5000 has been additionally processed using Pan-STARRS DR3 $i-$band data.
Figure 22: The distribution of ${r}_{\mathrm{ABmag}}-{i}_{\mathrm{ABmag}}$ for radio galaxies commonly in catalogs using UNIONS2000 (cyan) and UNIONS5000 (blue).
Appendix F Catalog columns
We show delivered columns of our UNIONS-VLASS radio galaxy catalog in Table LABEL:tab:_catalog_columns, including VLASS whose radio source is set as the reference at 3 GHz, UNIONS (optical; ugriz), SDSS (optical; ugriz), Pan-STARRS (optical to NIR; grizy), AllWISE (NIR to MIR; 3.4, 4.6, 12, and 22 $µm\micron$m), unWISE (NIR; 3.4 and 4.6 $µm\micron$m), FIRST (1.4 GHz), and LoTSS (144 MHz), as well as properties calculated in this work.
Table 3: Columns for the Main Catalog
Column name Type Units Description
Columns for VLASS (Gordon et al., 2021)
Component_name string Unique name of the PyBDSF component
RA double deg Right ascension of component [deg]
DEC double deg Declination of component [deg]
E_RA double deg Error in RA [deg]
E_DEC double deg Error in dec [deg]
Total_flux double mJy Integrated flux of the component [mJy]
E_Total_flux double mJy Error in integrated flux [mJy]
Peak_flux double mJy/beam Peak flux of component [mJy/beam]
E_Peak_flux double mJy/beam Error in component peak flux [mJy/beam]
Maj double arcsec Component major axis size (FWHM) [arcsec]
E_Maj double arcsec Error in Maj [arcsec]
Min double arcsec Component minor axis size (FWHM) [arcsec]
E_Min double arcsec Error in Min [arcsec]
PA double deg Component position angle east of north [deg]
E_PA double deg Error in PA [deg]
S_Code string PyBDSF source type
DC_Maj double arcsec Deconvolved component major axis size (FWHM) [arcsec]
E_DC_Maj double arcsec 1 sigma error in DC_Maj [arcsec]
DC_Min double arcsec Deconvolved component minor axis size (FWHM) [arcsec]
E_DC_Min double arcsec 1 sigma error in DC_Min [arcsec]
DC_PA double deg Deconvolved component position angle east of north [deg]
E_DC_PA double deg 1 sigma error in DC_PA [deg]
Peak_to_ring double Ratio of the Peak_flux to the maximum flux in annulus of $r=5\arcsec$ to $R=10\arcsec$
centred on component RA Dec
Duplicate_flag double Flag to denote the duplicate status of a component
Quality_flag double QA flag to deal with spurious detections and duplicates due to
overlap between VLASS tile edges
Source_name string Name of the distinct radio source to which this component has been assigned
Source_type string The type of source the component has been assigned to
NN_dist double arcsec Angular distance to nearest other component in the catalogue [arcsec]
BMAJ double arcsec Major axis of the beam [arcsec]
BMIN double arcsec Minor axis of the beam [arcsec]
BPA double deg Beam position angle east of north [deg]
Columns for UNIONS (Gwyn et al., 2025)
ALPHA_J2000 Double deg
DELTA_J2000 Double deg
MAG_GAAP_u Float mag u-band GAaP magnitude optimal min_aper
MAGERR_GAAP_u Float mag u-band GAaP magnitude error optimal min_aper
FLAG_GAAP_u Short GAaP Flag for MAG_GAAP_u optimal min_aper
MAG_GAAP_g Float mag g-band GAaP magnitude optimal min_aper
MAGERR_GAAP_g Float mag g-band GAaP magnitude error optimal min_aper
FLAG_GAAP_g Short GAaP Flag for MAG_GAAP_g optimal min_aper
MAG_GAAP_r Float mag r-band GAaP magnitude optimal min_aper
MAGERR_GAAP_r Float mag r-band GAaP magnitude error optimal min_aper
FLAG_GAAP_r Short GAaP Flag for MAG_GAAP_r optimal min_aper
MAG_GAAP_i Float mag i-band GAaP magnitude optimal min_aper
MAGERR_GAAP_i Float mag i-band GAaP magnitude error optimal min_aper
FLAG_GAAP_i Short GAaP Flag for MAG_GAAP_i optimal min_aper
MAG_GAAP_z Float mag z-band GAaP magnitude optimal min_aper
MAGERR_GAAP_z Float mag z-band GAaP magnitude error optimal min_aper
FLAG_GAAP_z Short GAaP Flag for MAG_GAAP_z optimal min_aper
Z_B Double 9-band BPZ redshift estimate; peak of posterior probability distribu
Z_B_MIN Double Lower bound of the 68% confidence interval of Z_B
Z_B_MAX Double Upper bound of the 68% confidence interval of Z_B
Columns for SDSS DR16 (Ahumada et al., 2020)
objID Long SDSS unique object identifier (1)
RA_ICRS Double deg Right Ascension of the object (ICRS) at Epoch="MJD" (ra)
DE_ICRS Double deg Declination of the object (ICRS) at Epoch="MJD" (dec)
mode Short [1/4] PhotoMode (1=primary, 2=secondary, 3=family, 4=outside)
class String [0/8] Type of object (3=galaxy, 6=star) (type) (2)
clean Short [0/1] Clean photometry flag (1=clean; 0=unclean)
e_RA_ICRS Double arcsec [0/51703] Mean error on RAdeg (raErr)
e_DE_ICRS Double arcsec [0/22795] Mean error on DEdeg (decErr)
umag Float mag [4/37.3] Model magnitude in u filter AB scale (u) (7)
gmag Float mag [3.7/39.6] Model magnitude in g filter AB scale (g) (7)
rmag Float mag [3.8/40.4] Model magnitude in r filter AB scale (r) (7)
imag Float mag [2.6/39.5] Model magnitude in i filter AB scale (i) (7)
zmag Float mag [3.5/37.2] Model magnitude in z filter AB scale (z) (7)
e_umag Double mag [8.9e-8/3444] Mean error on umag (err_u)
e_gmag Double mag [3.8e-8/5955] Mean error on gmag (err_g)
e_rmag Double mag [5.2e-8/915] Mean error on rmag (err_r)
e_imag Double mag [0/634] Mean error on imag (err_i)
e_zmag Double mag [9.4e-8/3529] Mean error on zmag (err_z)
zsp Double [-0.012/7.06] Spectroscopic final redshift (when SpObjID>0) (spec_z) (6)
e_zsp Double [-6/30689] Mean error on zsp (negative for bad fit) (spec_zErr) (6)
f_zsp Short [0/902] Zwarning flag (spec_zWarning) (6)(9)
zph Float [5e-4/1] Photometric redshift estimated by robust fit to nearest neighbors
in a reference set (photoz_z) (12)
e_zph Float [0.003/0.3] Estimated error of the photometric redshift (photoz_zErr) (12)
$\langle\mathrm{zph}\rangle$ Float [0.009/0.9] Average redshift of the nearest neighbors; if significantly different
from zph this might be a better estimate than zph (photoz_nnAvgZ) (12)
Q Short [1/3] Quality of the observation:
1$=$bad; 2$=$barely acceptable; 3$=$good (field_quality) (13)
SDSS16 String SDSS-DR16 name based on J2000 position; provided by CDS
Sp-ID String Spectroscopic Plate-MJD-Fiber identifier (plate-mjd-fiberID)
Columns for Pan-STARRS DR1 (Chambers et al., 2016)
RAJ2000 Double deg Right ascension (J2000) (weighted mean) at epoch "Epoch" (raMean) (1)
DEJ2000 Double deg Declination (J2000) (weighted mean) at epoch "Epoch" (decMean) (1)
objID Long Unique object identifier (objID)
f_objID Long Information flag bitmask indicating details of the photometry (objInfoFlag) (2)
Qual Short Subset of objInfoFlag denoting whether this object is real
or a likely false positive (qualityFlag) (3)
e_RAJ2000 Double arcsec Right ascension standard deviation from single epoch detections (raMeanErr)
e_DEJ2000 Double arcsec Declination standard deviation from single epoch detections (decMeanErr)
Epoch Double d Modified Julian Date (MJD) of the mean epoch (epochMean)
Ns Short [0/50] Number of stack detections (nStackDetections)
Nd Short [3/369] Number of single epoch detections in all filters (nDetections)
gmag Double mag [-33/33] Mean PSF AB magnitude from g filter (4866AA) detections (gMeanPSFMag)
e_gmag Float mag [0/0.5] Error in gmag (gMeanPSFMagErr)
gKmag Double mag Mean Kron (1980) AB magnitude from g filter detections (gMeanKronMag)
e_gKmag Float mag Error in gKmag (gMeanKronMagErr)
gFlags Integer Information flag bitmask for mean object from g filter detections (gFlags) (4)
rmag Double mag [-6/30] Mean PSF AB magnitude from r filter (6215AA) detections (rMeanPSFMag)
e_rmag Float mag [0/0.5] Error in rmag (rMeanPSFMagErr)
rKmag Double mag Mean Kron (1980) AB magnitude from r filter detections (rMeanKronMag)
e_rKmag Float mag Error in rKmag (rMeanKronMagErr)
rFlags Integer Information flag bitmask for mean object from r filter detections (rFlags) (4)
imag Double mag [-5/30] Mean PSF AB magnitude from i filter (7545AA) detections (iMeanPSFMag)
e_imag Float mag [0/0.5] Error in imag (iMeanPSFMagErr)
iKmag Double mag Mean Kron (1980) AB magnitude from i filter detections (iMeanKronMag)
e_iKmag Float mag Error in iKmag (iMeanKronMagErr)
iFlags Integer Information flag bitmask for mean object from i filter detections (iFlags) (4)
zmag Double mag [-5/29] Mean PSF AB magnitude from z filter (8679AA) detections (zMeanPSFMag)
e_zmag Float mag [0/0.5] Error in zmag (zMeanPSFMagErr)
zKmag Double mag Mean Kron (1980) AB magnitude from z filter detections (zMeanKronMag)
e_zKmag Float mag Error in zKmag (zMeanKronMagErr)
zFlags Integer Information flag bitmask for mean object from z filter detections (zFlags) (4)
ymag Double mag [-5/29] Mean PSF AB magnitude from y filter (9633AA) detections (yMeanPSFMag)
e_ymag Float mag [0/0.5] Error in ymag (yMeanPSFMagErr)
yKmag Double mag Mean Kron (1980) AB magnitude from y filter detections (yMeanKronMag)
e_yKmag Float mag Error in yKmag (yMeanKronMagErr)
yFlags Integer Information flag bitmask for mean object from y filter detections (yFlags) (4)
Columns for AllWISE (Cutri et al., 2021)
AllWISE String WISE All-Sky Release Catalog name, based on J2000 position, (designation)
RAJ2000 Double deg Right ascension (J2000)
DEJ2000 Double deg Declination (J2000)
Im String Image of the 4 bands from the IRSA server (IPAC)
w1mpro Float mag W1 magnitude (3.35um)
w1sigmpro Float mag Mean error on W1 magnitude
w2mpro Float mag W2 magnitude (4.6um)
w2sigmpro Float mag Mean error on W2 magnitude
w3mpro Float mag W3 magnitude (11.6um)
w3sigmpro Float mag Mean error on W3 magnitude
w4mpro Float mag W4 magnitude (22.1um)
w4sigmpro Float mag Mean error on W4 magnitude
Jmag Float mag 2MASS J magnitude (1.25um)
e_Jmag Float mag Mean error on J magnitude
Hmag Float mag 2MASS H magnitude (1.65um)
e_Hmag Float mag Mean error on H magnitude
Kmag Float mag 2MASS Ks magnitude (2.17um)
e_Kmag Float mag Mean error on Ks magnitude
cc_flags String [0DHOPdhop] Contamination and confusion flag, one per band (cc_flags) (2)
ext_flg Short [0-5] Extended source flag (ext_flg) (3)
var_flg String Variability flag, one per band (var_flg) (4)
ph_qual String [ABCUXZ] Photometric quality flag (8)
Columns for unWISE (Schlafly et al., 2019)
unwise_objid char unique object ID
ra double deg right ascension: W1 position, if available; otherwise W2 position
dec double deg declination: W1 position, if available; otherwise W2 position
flux_1 double ’nMgy’ flux, band 1
dflux_1 double ’nMgy’ statistical uncertainty in flux_1
flux_2 double ’nMgy’ flux, band 2
dflux_2 double ’nMgy’ statistical uncertainty in flux_2
fluxlbs_1 double ’nMgy’ local-background-subtracted flux, band 1
dfluxlbs_1 double ’nMgy’ uncertainty in fluxlbs_1
fluxlbs_2 double ’nMgy’ local-background-subtracted flux, band 2
dfluxlbs_2 double ’nMgy’ uncertainty in fluxlbs_2
qf_1 double quality factor, band 1: PSF-weighted fraction of pixels contributing to this detection
qf_2 double quality factor, band 2: PSF-weighted fraction of pixels contributing to this detection
rchi2_1 double average $\mathrm{\chi^{2}}$ per pixel, weighted by PSF, band 1
rchi2_2 double average $\mathrm{\chi^{2}}$ per pixel, weighted by PSF, band 2
fracflux_1 double fraction of flux in this object’s PSF that comes from this object, band 1
fracflux_2 double fraction of flux in this object’s PSF that comes from this object, band 2
spread_model_1 double SExtractor spread_model parameter, band 1
dspread_model_1 double uncertainty in spread_model_1
spread_model_2 double SExtractor spread_model parameter, band 2
dspread_model_2 double uncertainty in spread_model_2
fwhm_1 double pix full-width at half-maximum of PSF, band 1
fwhm_2 double pix full-width at half-maximum of PSF, band 2
Columns for FIRST (Helfand et al., 2015)
FIRST String FIRST Source designation
RAJ2000 String Right Ascension J2000
DEJ2000 String Declination J2000
p(S) Float [0,1] Probability of being a side lobe
Fpeak Double mJy Peak flux density at 1.4GHz
Fint Double mJy Integrated flux density at 1.4GHz
Rms Float mJy Local noise estimate
Maj Float arcsec Major axis (FWHM)
Min Float arcsec Minor axis (FWHM)
PA Float deg [0/180] Position angle
N1 String [0/10] Number of SDSS-DR10 counterparts
c1 Character [sg-] SDSS class: s=star, g=galaxy
N2 String Number of 2MASS counterparts
Columns for LoTSS DR2 (Shimwell et al., 2022)
Source String The radio name of the source, automatically generated from RA and DEC
RAJ2000 Double deg Right ascension (J2000) (RA)
DEJ2000 Double deg Declination (J2000) (DEC)
Speak Double mJy/beam The peak Stokes I flux density per beam of the source (Peak_flux)
SpeakTot Double mJy The total, integrated Stokes I flux density of the source
at the reference frequency (Total_flux)
Maj Float arcsec FWHM of the major axis of the source, INCLUDING convolution
with the 6-arcsec LOFAR beam (Maj)
Min Float arcsec FWHM of the minor axis of the source, INCLUDING convolution
with the 6-arcsec LOFAR beam (Min)
DCMaj Float arcsec The FWHM of the major axis of the source, after de-convolution
with the 6-arcsec LOFAR beam (DC_Maj)
DCMin Float arcsec The FWHM of the minor axis of the source, after de-convolution
with the 6-arcsec LOFAR beam (DC_Min)
PA Float deg The position angle of the major axis of the source measured east of north,
after de-convolution with the 6-arcsec LOFAR beam (PA)
DCPA Float deg The position angle of the major axis of the source measured east of north,
after de-convolution with the 6-arcsec LOFAR beam (DC_PA)
SCode String A code that defines the source structure
in terms of the fitted Gaussian components
MaskFract Float The fraction of the source that is in the CLEAN mask (Masked_Fraction)
Columns calculated in this work
flag_multi_opt int 0=VLASS source has unique UNIONS counterpart within the search radius;
1=VLASS source has multiple UNIONS counterparts within the search radius;
sep_guc_vlass float arcsec The angular separation between UNIONS and VLASS coordinates
sep_guc_first float arcsec The angular separation between UNIONS and FIRST coordinates
sep_guc_lotss float arcsec The angular separation between UNIONS and LoTSS coordinates
sep_vlass_first float arcsec The angular separation between VLASS and FIRST coordinates
sep_vlass_lotss float arcsec The angular separation between VLASS and LoTSS coordinates
sep_first_lotss float arcsec The angular separation between FIRST and LoTSS coordinates
alpha_144_1400 float The spectral index between LoTSS 144 MHz and FIRST 1.4 GHz
alpha_144_1400_err float Uncertainties of alpha_144_1400
alpha_1400_3000 float The spectral index between FIRST 1.4 GHz and VLASS 3 GHz
alpha_1400_3000_err float Uncertainties of alpha_1400_3000
alpha_144_3000 float The spectral index between LoTSS 144 MHz and VLASS 3 GHz
alpha_144_3000_err float Uncertainties of alpha_144_3000
r_obs float Observed radio loudness defined as the ratio of 1.4 GHz to $i$-band flux densities
luminosity_1p4 float W/Hz Spectral luminosity at 1.4 GHz, VLASS 3 GHz scaled if no FIRST detection
luminosity_3p0 float W/Hz Spectral luminosity at 3 GHz, alpha=-0.7 used if no FIRST detection
mag_vlass float mag Radio AB magnitude for VLASS
mag_first float mag Radio AB magnitude for FIRST
ls_vlass float kpc VLASS projected linear size calculated from the deconvolved major axis
ls_lotss float kpc LoTSS projected linear size calculated from the deconvolved major axis
flag_z_photo int -1$=$no valid photo-$z$;
[…]
Z_ref float preferred redshifts; Z_ref = SDSS spec-$z$ if available; otherwise Z_ref = Z_B_ref
extended int 1=Extended; 0=Not Extended. Classified following the criterion in § 3.6.3